### v3

# IMPLICATIVE BE ALGEBRAS AS ORTHOGONALITY SPACES

Abstract.
We obtain an orthogonality space by endowing an implicative involutive BE algebra with an appropriate orthogonality relation; for these spaces, we also study the particular case of implicative-orthomodular lattices. Moreover, we define the Sasaki projections as well as the commutativity relation on these structures, proving that an implicative involutive BE algebra is an implicative-orthomodular lattice if and only if the commutativity relation is symmetric. We also give characterizations of implicative-Boolean algebras, and we show that the center of an implicative-orthomodular lattice is an implicative-Boolean algebra. We prove that an implicative involutive BE algebra is an implicative-orthomodular lattice if and only if it admits a full Sasaki set of projections. Finally, based on Sasaki maps on implicative involutive BE algebras, we introduce the notion of Sasaki spaces, proving that, in the case when a complete implicative involutive BE algebra admits a full Sasaki set of projections, it is a Sasaki space. We also characterize the Dacey spaces defined by implicative involutiv BE algebras.
Keywords: orthogonality space, implicative involutive BE algebra, implicative-orthomodular lattice, implicative-Boolean algebra, Sasaki projection, commutativity relation, Sasaki space, Dacey space
AMS Subject Classification (2020): 06C15, 03G25, 06A06, 81P10
1. Introduction
A mathematical description of the structure of random events of a quantum mechanical system is one of the most important problem formulated by G. Birkhoff and J. von Neumann in their famous paper “The Logic of Quantum Mechanics” ([2]). The behavior of physical systems which describe the events concerning elementary particles is different from that of physical systems observed in classical physics, and another approach had to be developed. The first model proposed by G. Birkhoff and J. von Neumann was based on Hilbert spaces and their operators, taking into consideration that in quantum mechanics the state of a physical system is represented by a vector in a Hilbert space. A fundamental role in the study of Hilbert spaces is played by orthomodular posets and orthomodular lattices, and the developing of theories of orthomodular structures is still a central topic of research.
Recall that (see for example [16]) an orthogonality space (or an orthoset) $(X,\perp)$ is a set $X$ equipped with a symmetric and irreflexive binary relation $\perp$, called the orthogonality relation. We call $x,y\in X$ orthogonal elements if $x\perp y$. If $A\subseteq X$, denote $A^{\perp}=\{x\in X\mid x\perp a$, for all $a\in A\}$. The subset $A$ is called orthoclosed if $A=A^{\perp\perp}$, and denote by $\mathcal{C}(X,\perp)$ the set of all orthoclosed subsets of $X$. J.R. Dacey was the first to study the orthogonality spaces from the point of view of the foundations of quantum physics ([6]; see also [19]). Remarkable results on orthogonality spaces were recently presented in [16, 18, 19, 22, 23]. Orthomodular lattices were first introduced by G. Birkhoff and J. von Neumann ([2]), and, independently, by K. Husimi ([10]) by studying the structure of the lattice of projection operators on a Hilbert space. They introduced the orthomodular lattices as the event structures describing quantum mechanical experiments. Later, orthomodular lattices and orthomodular posets were considered as the standard quantum logic [21] (see also [7]). Complete studies on orthomodular lattices are presented in the monographs [1, 14]. An orthomodular lattice (OML for short) is an ortholattice $(X,\wedge,\vee,^{\prime},0,1)$ verifying:
$(OM)$ $(x\wedge y)\vee((x\wedge y)^{\prime}\wedge x)=x$, for all $x,y\in X$, or, equivalently
$(OM^{\prime})$ $x\vee(x^{\prime}\wedge y)=y$, whenever $x\leq y$ (where $x\leq y$ iff $x=x\wedge y$) ([17]).
The orthomodular lattices were redefined and studied by A. Iorgulescu ([12]) based on involutive m-BE algebras, and she gave two generalizations of orthomodular lattices: orthomodular softlattices and orthomodular widelattices. Based on implicative involutive BE algebras, we redefined in [4] the orthomodular lattices, by introducing and studying the implicative-orthomodular lattices (IOML for short), and we gave certain characterizations of these structures. The distributivity of implicative-orthomodular lattices was investigated in [5]. The notion of Sasaki projections on orthocomplemented lattices was introduced by U. Sasaki ([20]), and this notion was extended to other structures ([8, 9, 3]).
In this paper, we obtain an orthogonality space by endowing an implicative involutive BE algebra with an appropriate orthogonality relation, and we investigate the properties of these spaces. Special results are proved for the particular case of implicative-orthomodular lattices as orthogonality spaces. We define and study the Sasaki projections, and based on these notions we define and characterize the commutativity relation on implicative-orthomodular lattices. Moreover, we prove that an implicative involutive BE algebra is an implicative-orthomodular lattice if and only if the commutativity relation is symmetric. Based on the divisibility condition, we introduce and characterize the implicative-Boolean algebras, proving that the center of an implicative-orthomodular lattice is an implicative-Boolean algebra. We also introduce the notion of Sasaki set of projections, proving that an implicative involutive BE algebra is an implicative-orthomodular lattice if and only if it admits a full Sasaki set of projections. Inspired from [16], we define the Sasaki maps on the associated orthogonality space of an implicative involutive BE algebras and we show that, in the case of implicative-orthomodular lattices, the restriction and corestriction of a Sasaki projection is a Sasaki map. We also characterize the Dacey spaces defined by implicative involutiv BE algebras. Finally, we prove that, in the case when a complete implicative involutive BE algebra admits a full Sasaki set of projections, then it is a Sasaki space.
2. Implicative involutive BE algebras as orthogonality spaces
In this section, given an implicative involutive BE algebra $(X,\rightarrow,1)$, we define an orthogonality relation $\perp$ on $X^{\prime}=X\setminus\{0\}$, and we prove that $(X^{\prime},\perp)$ is an orthogonality space. We recall from [4] the notions of implicative-orthomodular lattices and their properties, proving new properties and a new characterization theorem.
BE algebras were introduced in [15] as algebras $(X,\rightarrow,1)$ of type $(2,0)$ satisfying the following conditions, for all $x,y,z\in X$: $(BE_{1})$ $x\rightarrow x=1;$ $(BE_{2})$ $x\rightarrow 1=1;$ $(BE_{3})$ $1\rightarrow x=x;$ $(BE_{4})$ $x\rightarrow(y\rightarrow z)=y\rightarrow(x\rightarrow z)$. A relation $\leq$ is defined on $X$ by $x\leq y$ iff $x\rightarrow y=1$. A BE algebra $X$ is bounded if there exists $0\in X$ such that $0\leq x$, for all $x\in X$. In a bounded BE algebra $(X,\rightarrow,0,1)$ we define $x^{*}=x\rightarrow 0$, for all $x\in X$. A bounded BE algebra $X$ is called involutive if $x^{**}=x$, for any $x\in X$.
Lemma 2.1.
[…]
$(1)$ $x\rightarrow(y\rightarrow x)=1;$ $(2)$ $x\leq(x\rightarrow y)\rightarrow y$.
If $X$ is bounded, then:
$(3)$ $x\rightarrow y^{*}=y\rightarrow x^{*};$ $(4)$ $x\leq x^{**}$.
If $X$ is involutive, then:
$(5)$ $x^{*}\rightarrow y=y^{*}\rightarrow x;$ $(6)$ $x^{*}\rightarrow y^{*}=y\rightarrow x;$
$(7)$ $(x\rightarrow y)^{*}\rightarrow z=x\rightarrow(y^{*}\rightarrow z);$ $(8)$ $x\rightarrow(y\rightarrow z)=(x\rightarrow y^{*})^{*}\rightarrow z;$
$(9)$ $(x^{*}\rightarrow y)^{*}\rightarrow(x^{*}\rightarrow y)=(x^{*}\rightarrow x)^{*}\rightarrow(y^{*}\rightarrow y)$.
In a BE algebra $X$, we define the additional operation:
                           $x\Cup y=(x\rightarrow y)\rightarrow y$.
If $X$ is involutive, we define the operation:
                           $x\Cap y=((x^{*}\rightarrow y^{*})\rightarrow y^{*})^{*}$,
and the relations $\leq_{Q}$ and $\leq_{L}$ by:
                           $x\leq_{Q}y$ iff $x=x\Cap y$ and $x\leq_{L}y$ iff $x=(x\rightarrow y^{*})^{*}$.
Obviously, $x\leq_{L}y$ iff $x^{*}=x\rightarrow y^{*}=y\rightarrow x^{*}$.
Proposition 2.2.
Let $X$ be an involutive BE algebra. The following hold, for all $x,y,z,u\in X$:
$(1)$ $x\leq_{Q}y$ implies $x=y\Cap x$ and $y=x\Cup y;$
$(2)$ $\leq_{Q}$ is reflexive and antisymmetric;
$(3)$ $x\Cap y=(x^{*}\Cup y^{*})^{*}$ and $x\Cup y=(x^{*}\Cap y^{*})^{*};$
$(4)$ $x\leq_{Q}y$ implies $x\leq y;$
$(5)$ $x,y\leq_{Q}z$ and $z\rightarrow x=z\rightarrow y$ imply $x=y$ (cancellation law);
$(6)$ $x\leq_{L}y$ implies $x\leq y;$
$(7)$ $\leq_{L}$ is an order relation on $X;$
$(8)$ $z\leq_{L}x$ and $z\leq_{L}y$ imply $z\leq_{L}(x\rightarrow y^{*})^{*};$
$(9)$ $(x\rightarrow y^{*})^{*}\rightarrow(z\rightarrow u^{*})=(x\rightarrow z^{*})^{*}\rightarrow(y\rightarrow u^{*})$.
Proof.
$(1)$-$(7)$ See [4].
$(8)$ Since $z\leq_{L}x,y$ implies $z=(z\rightarrow x^{*})^{*}=(z\rightarrow y^{*})^{*}$, we get $(z\rightarrow(x\rightarrow y^{*}))^{*}=(x\rightarrow(z\rightarrow y^{*}))^{*}=(x\rightarrow z^{*})^{*}=(z\rightarrow x^{*})^{*}=z$, hence $z\leq_{L}(x\rightarrow y^{*})^{*}$.
$(9)$ Using Lemma 2.1$(7)$, we have:
                   $(x\rightarrow y^{*})^{*}\rightarrow(z\rightarrow u^{*})=x\rightarrow(y\rightarrow(z\rightarrow u^{*}))$
                                                     $=x\rightarrow(z\rightarrow(y\rightarrow u^{*}))$
                                                     $=(x\rightarrow z^{*})^{*}\rightarrow(y\rightarrow u^{*})$. ∎
The implicative involutive BE algebras were introduced by A. Iorgulescu in [13], and their properties were also studied in [4].
Definition 2.3.
A BE algebra is called implicative if it satisfies the following condition, for all $x,y\in X$:
$(Impl)$ $(x\rightarrow y)\rightarrow x=x$.
Lemma 2.4.
([4]) Let $(X,\rightarrow,^{*},1)$ be an implicative involutive BE algebra. Then $X$ verifies the following axioms, for all $x,y\in X$:
$(iG)$ $x^{*}\rightarrow x=x$, or equivalently, $x\rightarrow x^{*}=x^{*};$
$(Iabs$-$i)$ $(x\rightarrow(x\rightarrow y))\rightarrow x=x;$
$(Pimpl)$ $x\rightarrow(x\rightarrow y)=x\rightarrow y$.
Lemma 2.5.
([4]) Let $(X,\rightarrow,^{*},1)$ be an involutive BE algebra. The following are equivalent:
$(a)$ $X$ is implicative;
[…]
Lemma 2.6.
Let $X$ be an implicative involutive BE algebra. The following hold, for all $x,y,z,u\in X$:
$(1)$ $x\leq_{L}y$ iff $y^{*}\leq_{L}x^{*};$ $(2)$ $x\leq_{Q}y$ implies $x\leq_{L}y;$
$(3)$ $x\leq_{L}y\rightarrow x$ and $x\leq_{L}x^{*}\rightarrow y;$ $(4)$ $(x\rightarrow y^{*})^{*}\leq_{L}x,y;$
$(5)$ $x^{*}=x\rightarrow y$ iff $y^{*}=y\rightarrow x;$ $(6)$ $x\Cap y=x$ implies $x\Cap y^{*}=0;$
$(7)$ $x\leq_{L}y$ implies $x\Cap y^{*}=0;$
If $x\leq_{L}y$, then:
$(8)$ $y\rightarrow z\leq_{L}x\rightarrow z$ and $z\rightarrow x\leq_{L}z\rightarrow y;$
$(9)$ $x\Cup z\leq_{L}y\Cup z$ and $x\Cap z\leq_{L}y\Cap z;$
$(10)$ if $x,y\leq_{L}z$, then $x^{*}\rightarrow y\leq_{L}z;$
$(11)$ $(x\rightarrow y^{*})\rightarrow(x\rightarrow y)^{*}\leq_{L}x;$
$(12)$ if $x\leq_{L}y$ and $z\leq_{L}u$, then $x^{*}\rightarrow z\leq_{L}y^{*}\rightarrow u$.
Proof.
$(1)$-$(4)$ See [4].
$(5)$ If $x^{*}=x\rightarrow y$, using $(Impl)$ we get: $y\rightarrow x=y\rightarrow(x\rightarrow y)^{*}=(x\rightarrow y)\rightarrow y^{*}=(y^{*}\rightarrow x^{*})\rightarrow y^{*}=y^{*}$. Conversely, if $y^{*}=y\rightarrow x$, then $x\rightarrow y=x\rightarrow(y\rightarrow x)^{*}=(y\rightarrow x)\rightarrow x^{*}=(x^{*}\rightarrow y^{*})\rightarrow x^{*}=x^{*}$.
$(6)$ If $x\Cap y=x$, then $(x^{*}\rightarrow y^{*})\rightarrow y^{*}=x^{*}$, so that $x^{*}=y\rightarrow(y\rightarrow x)^{*}$. Using $(Impl)$, we get: $x\Cap y^{*}=((x^{*}\rightarrow y)\rightarrow y)^{*}=(((y\rightarrow(y\rightarrow x)^{*})\rightarrow y)\rightarrow y)^{*}=(y\rightarrow y)^{*}=0$.
$(7)$ Since $x\leq_{L}y$ implies $x^{*}=x\rightarrow y^{*}$, the proof is similar to $(6)$.
$(8)$ Since $x\leq_{L}y$, we have $x^{*}=x\rightarrow y^{*}$, so that
                  $x\rightarrow z=z^{*}\rightarrow x^{*}=z^{*}\rightarrow(x\rightarrow y^{*})=x\rightarrow(z^{*}\rightarrow y^{*})=x\rightarrow(y\rightarrow z)$.
Applying $(Impl)$, it follows that:
          $((y\rightarrow z)\rightarrow(x\rightarrow z)^{*})^{*}=((x\rightarrow z)\rightarrow(y\rightarrow z)^{*})^{*}$
                                            $=((x\rightarrow(y\rightarrow z))\rightarrow(y\rightarrow z)^{*})^{*}$
                                            $=(((y\rightarrow z)^{*}\rightarrow x^{*})\rightarrow(y\rightarrow z)^{*})^{*}$
                                            $=(y\rightarrow z)^{**}=y\rightarrow z$,
hence $y\rightarrow z\leq_{L}x\rightarrow z$. Moreover, $x\leq_{L}y$ implies $y^{*}\leq_{L}x^{*}$, so that $x^{*}\rightarrow z^{*}\leq_{L}y^{*}\rightarrow z^{*}$, that is $z\rightarrow x\leq_{L}z\rightarrow y$.
$(9)$ From $x\leq_{L}y$, using $(8)$, we get $y\rightarrow z\leq_{L}x\rightarrow z$ and $(x\rightarrow z)\rightarrow z\leq_{L}(y\rightarrow z)\rightarrow z$, so that $x\Cup z\leq_{L}y\Cup z$. Moreover, $y^{*}\leq_{L}x^{*}$ implies $y^{*}\Cup z^{*}\leq_{L}x^{*}\Cup z^{*}$, hence $(x^{*}\Cup z^{*})^{*}\leq_{L}(y^{*}\Cup z^{*})^{*}$, that is $x\Cap z\leq y\Cap y\Cap z$.
$(10)$ $x,y\leq_{L}z$ implies $z^{*}\leq_{L}x^{*},y^{*}$, so that by $(8)$ and $(iG)$ we have: $x^{*}\rightarrow y\leq_{L}z^{*}\rightarrow y\leq_{L}z^{*}\rightarrow z=z$.
$(11)$ It follows by $(10)$, since $(x\rightarrow y^{*})^{*}\leq_{L}x$ and $(x\rightarrow y)^{*}\leq_{L}x$.
$(12)$ From $x\leq_{L}y$ we have $y^{*}\leq_{L}x^{*}$, hence $x^{*}\rightarrow z\leq_{L}y^{*}\rightarrow z\leq_{L}y^{*}\rightarrow u$. ∎
Definition 2.7.
Given an implicative involutive BE algebra $(X,\rightarrow,^{*},1)$, the elements $x,y\in X$ are called orthogonal, denoted by $x\perp y$, if $x^{*}=x\rightarrow y$.
Lemma 2.8.
Let $X$ be an implicative involutive BE algebra. The following hold, for all $x,y\in X$:
$(1)$ $x\perp y$ iff $y\perp x;$ $(2)$ $x\perp x$ iff $x=0;$ $(3)$ $0\perp x;$
$(4)$ $1\perp x$ iff $x=0;$ $(5)$ $x\leq_{L}y$ implies $x\perp y^{*}$.
Proof.
$(1)$ follows by Lemma 2.6$(5)$, while $(2)$-$(5)$ are straightforward. ∎
It follows that $\perp$ is an orthogonality relation on $X^{\prime}=X\setminus\{0\}$, so that $(X^{\prime},\perp)$ is an orthogonality space. If $A\subseteq X^{\prime}$, denote by $A^{\perp}=\{x\in X^{\prime}\mid x\perp a$, for all $a\in A\}$. For $A=\{x\}$ ($A$ is a singleton), we denote $x^{\perp}$ instead of $\{x\}^{\perp}$.
Lemma 2.9.
Let $X$ be an involutive BE algebra. The following are equivalent:
$(IOM)$ $x\Cap(y\rightarrow x)=x$, for all $x,y\in X;$
$(IOM^{\prime})$ $x\Cap(x^{*}\rightarrow y)=x$, for all $x,y\in X;$
$(IOM^{\prime\prime})$ $x\Cup(x\rightarrow y)^{*}=x$, for all $x,y\in X$.
Proof.
The proof is straightforward. ∎
Definition 2.10.
An implicative-orthomodular lattice (IOML for short) is an implicative involutive BE algebra satisfying any one of the equivalent conditions from Lemma 2.9.
Remark 2.11.
An implicative involutive BE algebra ia called an implicative-ortholattice ([11, Def. 3.29]). The implicative involutive BE algebras $(X,\rightarrow,^{*},1)$ are term-equivalent to ortholattices $(X,\wedge,\vee,^{\prime},0,1)$, by the mutually inverse transformations:
                           $\varphi:$  $x\rightarrow y:=(x\wedge y^{\prime})^{\prime}$   and   $\psi:$  $x\wedge y:=(x\rightarrow y^{*})^{*}$,
and the relation $x\vee y:=(x^{\prime}\wedge y^{\prime})^{\prime}=x^{*}\rightarrow y$. The unary operation “” is defined as $x^{*}:=x\rightarrow 0=(x\wedge 0^{\prime})^{\prime}=x^{\prime}$.
Also, by these transformations the orthomodular lattices are term-equivalent to implicative-orthomodular lattices.
We have the following characterization of implicative-orthomodular lattices.
Theorem 2.12.
([5]) Let $X$ be an implicative involutive BE algebra. The following are equivalent:
$(a)$ $X$ is an implicative-orthomodular lattice;
$(b)$ $\leq_{L}$ implies $\leq_{Q};$
$(c)$ $x\leq_{L}y$ implies $y=y\Cup x$.
Corollary 2.13.
If $X$ is an implicative-orthomodular lattice, then $\leq_{Q}\Longleftrightarrow\leq_{L}$.
Proof.
It follows by Theorem 2.12 and Lemma 2.6. ∎
Proposition 2.14.
([4]) Let $X$ be an implicative-orthomodular lattice. Then the following hold, for all $x,y,z\in X$:
$(1)$ $x\rightarrow(y\Cap x)=x\rightarrow y;$
$(2)$ $(x\Cup y)\rightarrow(x\rightarrow y)^{*}=y^{*};$
$(3)$ $x\Cap((y\rightarrow x)\Cap(z\rightarrow x))=x;$
$(4)$ $(x\rightarrow y)\rightarrow(y\Cap x)=x;$
$(5)$ $x\leq_{L}y$ and $x\leq y$ imply $x=y;$
$(6)$ $x\Cap y\leq_{L}y\leq_{L}x\Cup y;$
$(7)$ $(x\Cap y)\rightarrow(y\Cap x)=1;$
$(8)$ $(x\Cup y)\rightarrow(y\Cup x)=1;$
$(9)$ $(x\Cup y)\rightarrow y=x\rightarrow y;$
$(10)$ $(x\Cap y)\Cap(y\Cap z)=(x\Cap y)\Cap z$.
Proposition 2.15.
Let $X$ be an implicative-orthomodular lattice. The following hold, for all $x,y,z\in X$:
$(1)$ $x\leq_{L}y$, $x\leq_{L}z$ imply $x\leq_{L}y\Cap z;$
$(2)$ $x\leq_{L}y$ implies $(z\Cap y)\Cap x=z\Cap x;$
$(3)$ if $x\leq y$, $y\leq_{L}x$ imply $x=y;$
$(4)$ $y\leq_{L}x$, $z\leq_{L}x$ imply $y\Cup z\leq_{L}x;$
$(5)$ $x\rightarrow(x\Cap y)=x\rightarrow y;$
$(6)$ $x\Cap y^{*}=0$ implies $x\Cap y=x;$
$(7)$ if $x,y\in X^{\prime}=X\setminus\{0\}$, then $x\perp y$ iff $x\Cap y=0$.
Proof.
$(1)$ From $x\leq_{L}y$, $x\leq_{L}z$ we have $x\leq_{Q}y$, $x\leq_{Q}z$, hence $x=x\Cap y$ and $x=x\Cap z$. Applying Proposition 2.14$(10)$, we get: $x\Cap(y\Cap z)=(x\Cap y)\Cap(y\Cap z)=(x\Cap y)\Cap z=x\Cap z=x$, hence $x\leq_{Q}y\Cap z$, that is $x\leq_{L}y\Cap z$.
$(2)$ From $x\leq_{L}y$, we have $x=(x\rightarrow y^{*})^{*}$, and Applying Proposition 2.14$(1)$, it follows that:
           $(z\Cap y)\Cap x=(((z\Cap y)^{*}\rightarrow x^{*})\rightarrow x^{*})^{*}=(((z\Cap y)^{*}\rightarrow(x\rightarrow y^{*}))\rightarrow x^{*})^{*}$
                           $=((x\rightarrow((z\Cap y)^{*}\rightarrow y^{*}))\rightarrow x^{*})^{*}=((x\rightarrow(y\rightarrow(z\Cap y)))\rightarrow x^{*})^{*}$
                           $=((x\rightarrow(y\rightarrow z))\rightarrow x^{*})^{*}=((x\rightarrow(z^{*}\rightarrow y^{*}))\rightarrow x^{*})^{*}$
                           $=((z^{*}\rightarrow(x\rightarrow y^{*}))\rightarrow x^{*})^{*}=((z^{*}\rightarrow x^{*})\rightarrow x^{*})^{*}=z\Cap x$.
$(3)$ From $x\leq y$, we get $y\Cap x=((y^{*}\rightarrow x^{*})\rightarrow x^{*})^{*}=((x\rightarrow y)\rightarrow x^{*})^{*}=(1\rightarrow x^{*})^{*}=x$. On the other hand, since $y\leq_{L}x$, we have $y^{*}=y\rightarrow x^{*}$, and we get:
                  $y\Cap x=((y^{*}\rightarrow x^{*})\rightarrow x^{*})^{*}=(((y\rightarrow x^{*})\rightarrow x^{*})\rightarrow x^{*})^{*}$
                           $=((y\Cup x^{*})\rightarrow x^{*})^{*}=(x\rightarrow(y^{*}\Cap x))^{*}$
                           $=(x\rightarrow y^{*})^{*}=y$ (by Prop. 2.14$(1)$).
It follows that $x=y$.
$(4)$ From $y,z\leq_{L}x$, we have $x^{*}\leq_{L}y^{*},z^{*}$, and applying $(1)$, we get $x^{*}\leq_{L}y^{*}\Cap z^{*}=(y\Cup z)^{*}$, that is $y\Cup z\leq_{L}x$.
$(5)$ Since by $(Impl)$, $x=((x^{*}\rightarrow y)\rightarrow x^{*})^{*}=(x\rightarrow(x^{*}\rightarrow y)^{*})^{*}$, we get:
                  $x\rightarrow(x\Cap y)=x\rightarrow((x^{*}\rightarrow y^{*})\rightarrow y^{*})^{*}$
                                    $=(x\rightarrow(x^{*}\rightarrow y)^{*})^{*}\rightarrow((x^{*}\rightarrow y^{*})\rightarrow y^{*})^{*}$
                                    $=x\rightarrow((x^{*}\rightarrow y^{*})\rightarrow((x^{*}\rightarrow y^{*})\rightarrow y^{*})^{*})$ (Lemma 2.1$(7)$)
                                    $=x\rightarrow(((x^{*}\rightarrow y^{*})\rightarrow y^{*})\rightarrow(x^{*}\rightarrow y^{*})^{*})$ (Lemma 2.1$(6)$)
                                    $=x\rightarrow((y\rightarrow(x^{*}\rightarrow y^{*})^{*})\rightarrow(x^{*}\rightarrow y^{*})^{*})$ (Lemma 2.1$(6)$)
                                    $=x\rightarrow(y\Cup(x^{*}\rightarrow y^{*})^{*})$
                                    $=x\rightarrow(y^{*}\Cap(x^{*}\rightarrow y^{*}))^{*}$
                                    $=x\rightarrow(y^{*})^{*}$ (by $(IOM)$)
                                    $=x\rightarrow y$.
$(6)$ Assume that $x\Cap y^{*}=0$, hence $(x^{*}\rightarrow y)\rightarrow y=1$, so that $x^{*}\rightarrow y\leq y$. On the other hand, $y\leq_{L}x^{*}\rightarrow y$, and by $(3)$, we get $y=x^{*}\rightarrow y=y^{*}\rightarrow x$. It follows that:
                  $x\Cap y=((x^{*}\rightarrow y^{*})\rightarrow y^{*})^{*}=((y\rightarrow x)\rightarrow y^{*})^{*}=((x^{*}\rightarrow y)\rightarrow x)\rightarrow y^{*})^{*}$
                           $=(((y^{*}\rightarrow x)\rightarrow x)\rightarrow y^{*})^{*}=((y^{*}\Cup x)\rightarrow y^{*})^{*}$
                           $=(y\rightarrow(y\Cap x^{*}))^{*}=(y\rightarrow x^{*})^{*}$ (by $(5)$).
                           $=((x^{*}\rightarrow y)\rightarrow x^{*})^{*}=(x^{*})^{*}=x$ (by $(Impl)$).
$(7)$ Assume that $x\perp y$, so that $x^{*}=x\rightarrow y$, and by $(Impl)$ we get:
                  $x\Cap y=((x^{*}\rightarrow y^{*})\rightarrow y^{*})^{*}=(((x\rightarrow y)\rightarrow y^{*})\rightarrow y^{*})^{*}$
                           $=(((y^{*}\rightarrow x^{*})\rightarrow y^{*})\rightarrow y^{*})^{*}=(y^{*}\rightarrow y^{*})^{*}=0$.
Conversely, if $x\Cap y=0$, then $(x^{*}\rightarrow y^{*})\rightarrow y^{*}=1$, so that $x^{*}\rightarrow y^{*}\leq y^{*}$. On the other hand, $y^{*}\leq_{L}x^{*}\rightarrow y^{*}$, and by $(3)$, we get $y^{*}=x^{*}\rightarrow y^{*}=y\rightarrow x$. Thus $y\perp x$, that is, $x\perp y$. ∎
Theorem 2.16.
Let $X$ be an implicative involutive BE algebra. The following are equivalent:
$(a)$ $X$ is an implicative-orthomodular lattice;
$(b)$ $(x\rightarrow y)\rightarrow(y\Cap x)=x$, for all $x,y\in X;$
$(c)$ $x\leq y$ and $y\leq_{L}x$ imply $x=y;$
$(d)$ $x\Cap y^{*}=0$ implies $x\Cap y=x;$
$(e)$ $x\rightarrow(x\Cap y)=x\rightarrow y$, for all $x,y\in X$.
Proof.
$(a)\Leftrightarrow(b)$ If $X$ is an IOML and $x,y\in X$, by Proposition 2.14$(4)$, $(x\rightarrow y)\rightarrow(y\Cap x)=x$. Conversely, let $X$ be an implicative involutive BE algebra such that $(x\rightarrow y)\rightarrow(y\Cap x)=x$, for all $x,y\in X$. Then we have:
                  $x\Cup(x\rightarrow y)^{*}=(x\rightarrow(x\rightarrow y)^{*})\rightarrow(x\rightarrow y)^{*}$
                                     $=((x\rightarrow y)\rightarrow x^{*})\rightarrow(x\rightarrow y)^{*}$
                                     $=((y^{*}\rightarrow x^{*})\rightarrow x^{*})\rightarrow(x\rightarrow y)^{*}$
                                     $=(y\Cap x)^{*}\rightarrow(x\rightarrow y)^{*}$
                                     $=(x\rightarrow y)\rightarrow(y\Cap x)=x$.
Hence $X$ satisfies condition $(IOM^{{}^{\prime\prime}})$, that is $X$ is an IOML.
$(a)\Leftrightarrow(c)$ Let $X$ be an IOML, and let $x,y\in X$ such that $x\leq y$ and $y\leq_{L}x$. Then by Proposition 2.15$(3)$, $x=y$. Conversely, let $X$ be an implicative involutive BE algebra satisfying condition $(c)$, and let $x,y\in X$ such that $x\leq_{L}y$. We have $(x\Cap y)\rightarrow x=x^{*}\rightarrow(x\Cap y)^{*}=x^{*}\rightarrow((x^{*}\rightarrow y^{*})\rightarrow y^{*})=(x^{*}\rightarrow y^{*})\rightarrow(x^{*}\rightarrow y^{*})=1$, hence $x\Cap y\leq x$. Since $x\leq_{L}y$, then $x=(x\rightarrow y^{*})^{*}$, so that $x^{*}=x\rightarrow y^{*}$. It follows that:
                   $(x\rightarrow(x\Cap y)^{*})^{*}=(x\rightarrow((x^{*}\rightarrow y^{*})\rightarrow y^{*}))^{*}$
                                          $=((x^{*}\rightarrow y^{*})\rightarrow(x\rightarrow y^{*}))^{*}$
                                          $=((x^{*}\rightarrow y^{*})\rightarrow x^{*})^{*}=(x^{*})^{*}=x$ (by $(Impl)$),
thus $x\leq_{L}x\Cap y$. By hypothesis, $x\Cap y\leq x$ and $x\leq_{L}x\Cap y$ implies $x=x\Cap y$, that is $x\leq_{Q}y$. We proved that $x\leq_{L}y$ implies $x\leq_{Q}y$, so that, by Theorem 2.12, it follows that $X$ is an IOML.
$(a)\Leftrightarrow(d)$ Let $X$ be an IOML, and let $x,y\in X$ such that $x\Cap y^{*}=0$. Then by Proposition 2.15$(6)$, $x\Cap y^{*}=0$ implies $x\Cap y=x$. Conversely, let $X$ be an implicative involutive BE algebra satisfying condition $(d)$. For all $x,y\in X$, using $(iG)$ we have:
                  $x\Cap(y\rightarrow x)^{*}=((x^{*}\rightarrow(y\rightarrow x))\rightarrow(y\rightarrow x))^{*}$
                                     $=((y\rightarrow(x^{*}\rightarrow x))\rightarrow(y\rightarrow x))^{*}$
                                     $=((y\rightarrow x)\rightarrow(y\rightarrow x))^{*}=0$.
By hypothesis, $x\Cap(y\rightarrow x)=x$, hence $X$ satisfies condition $(IOM)$, so that $X$ is an IOML.
$(a)\Leftrightarrow(e)$ If $X$ is an IOML, by Proposition 2.15$(5)$ condition $(e)$ is satisfied. Conversely, assume that $X$ is an implicative involutive BE algebra verifying condition $(e)$. It follows that $x\rightarrow(x\Cap(y\rightarrow x^{*}))=x\rightarrow(y\rightarrow x^{*})$, and using $(Impl)$ we get $x\rightarrow(x\rightarrow(y\rightarrow x^{*})^{*})^{*}=y\rightarrow x^{*}$, for all $x,y\in X$. Repacing $x$ by $x\rightarrow y$ and $y$ by $x^{*}$ and applying again $(Impl)$, we get $((x\rightarrow y)\rightarrow((x\rightarrow y)\rightarrow x^{*})^{*}=x$. Finally, $x\Cup(x\rightarrow y)^{*})=x$, for all $x,y\in X$, so that $X$ satisfies condition $(IOM^{\prime\prime})$. Hence $X$ is an IOML. ∎
Remarks 2.17.
Let $X$ be an implicative-orthomodular lattice, and let $x,y\in X$.
$(1)$ By Lemma 2.6$(6)$ and Proposition 2.15$(6)$, $x\Cap y=x$ iff $x\Cap y^{*}=0$, or equivalently, $x\leq_{Q}y$ iff $x\Cap y^{*}=0$.
$(2)$ By $(1)$ and Corollary 2.13, $x\leq_{Q}y$ iff $x\leq_{L}y$ iff $x\Cap y^{*}=0$.
3. Sasaki projections on implicative-orthomodular lattices
We introduce the notions of Sasaki projections and commutativity relations on implicative-orthomodular lattices, and we study their properties. We show that an implicative involutive BE algebra is an implicative-orthomodular lattice if and only if the commutativity relation is symmetric. Moreover, the composition of two Sasaki projections $\varphi_{a}$, $\varphi_{b}$ on an implicative-orthomodular lattice is commutative if and only if $a$ commutes with $b$.
Definition 3.1.
Let $X$ be an implicative involutive BE algebra, and let $a\in X$. A Sasaki projection on $X$ is a map $\varphi_{a}:X\longrightarrow X$ defined by $\varphi_{a}(x)=x\Cap a$.
Denote $\mathcal{SP}(X)=\{\varphi_{a}\mid a\in X\}$ the set of all Sasaki projections on $X$. We write $\varphi_{a}x$, $\varphi_{a}\varphi_{b}x$ and $\varphi^{*}_{a}x$, instead of $\varphi_{a}(x)$, $(\varphi_{a}\circ\varphi_{b})(x)$ and $(\varphi_{a}x)^{*}$, respectively.
Proposition 3.2.
Let $X$ be an implicative involutive BE algebra, and let $a\in X$. The following hold, for all $x\in X$:
$(1)$ $\varphi_{x}x=x$, $\varphi_{1}x=\varphi_{x}1=x$, $\varphi_{0}x=\varphi_{x}0=0$, $\varphi_{x}x^{*}=\varphi_{x^{*}}x=0;$
$(2)$ $a\leq_{L}x$ implies $\varphi_{a}x=a;$
$(3)$ $\varphi_{\varphi_{a}x}x=\varphi_{a}x;$
$(4)$ $a\leq_{Q}x$ implies $\varphi_{x}a=a;$
$(5)$ $\varphi_{a}$ is monotone.
Proof.
$(1)$, $(2)$ and $(4)$ are straightforward.
$(3)$ We have:
                  $\varphi_{\varphi_{a}x}x=((x^{*}\rightarrow\varphi^{*}_{a}x)\rightarrow\varphi^{*}_{a}x)^{*}=((x^{*}\rightarrow(x\Cap a)^{*})\rightarrow\varphi^{*}_{a}x)^{*}$
                             $=((x^{*}\rightarrow((x^{*}\rightarrow a^{*})\rightarrow a^{*}))\rightarrow\varphi^{*}_{a}x)^{*}$
                             $=(((x^{*}\rightarrow a^{*})\rightarrow(x^{*}\rightarrow a^{*}))\rightarrow\varphi^{*}_{a}x)^{*}$
                             $=(1\rightarrow\varphi^{*}_{a}x)^{*}=\varphi_{a}x$.
$(5)$ Let $x_{1},x_{2}\in X$ such that $x_{1}\leq_{L}x_{2}$. Since by Lemma 2.6$(9)$, $x_{1}\Cap a\leq_{L}x_{2}\Cap a$, we have $\varphi_{a}x_{1}\leq_{L}\varphi_{a}x_{2}$. Hence $\varphi_{a}$ is monotone. ∎
Proposition 3.3.
Let $X$ be an implicative-orthomodular lattice. The following hold, for all $a,b,x,y\in X$:
$(1)$ if $\varphi_{a}=\varphi_{b}=\varphi_{1}$, then $\varphi_{a\Cap b}=\varphi_{1}$. $(2)$ $\varphi_{a}\varphi_{a}x=\varphi_{a}x;$
$(3)$ $\varphi_{a}\varphi^{*}_{a}x=(a\rightarrow x)^{*};$ $(4)$ $\varphi_{a}\varphi^{*}_{a}x\leq_{L}x^{*};$
$(5)$ $\varphi_{a}x\leq_{L}y^{*}$ iff $\varphi_{a}y\leq_{L}x^{*};$ $(6)$ $\varphi_{a}\varphi_{b}a=\varphi_{a}b;$
$(7)$ if $a=a\Cap b$, then $\varphi_{a}=\varphi_{a}\varphi_{b}$.
Proof.
$(1)$ From $\varphi_{a}=\varphi_{b}=\varphi_{1}$, we have $x\Cap a=x\Cap b=x$, thus $x\leq_{Q}a,b$, so that $x\leq_{L}a,b$. Applying Proposition 2.15$(1)$, we get $x\leq_{L}a\Cap b$, hence $x\Cap(a\Cap b)=x$, for all $x\in X$. It follows that $\varphi_{a\Cap b}=\varphi_{1}$.
$(2)$ We have:
                  $\varphi_{a}\varphi_{a}x=(x\Cap a)\Cap a=(((x\Cap a)^{*}\rightarrow a^{*})\rightarrow a^{*})^{*}$
                             $=((a\rightarrow(x\Cap a))\rightarrow a^{*})^{*}=((a\rightarrow x)\rightarrow a^{*})^{*}$ (Prop. 2.14$(1)$)
                             $=((x^{*}\rightarrow a^{*})\rightarrow a^{*})^{*}=x\Cap a=\varphi_{a}x$.
$(3)$ Similarly we get:
                  $\varphi_{a}(\varphi_{a}x)^{*}=(x\Cap a)^{*}\Cap a=(((x\Cap a)\rightarrow a^{*})\rightarrow a^{*})^{*}$
                                $=((a\rightarrow(x\Cap a)^{*})\rightarrow a^{*})^{*}=((a\rightarrow((x^{*}\rightarrow a^{*})\rightarrow a^{*}))\rightarrow a^{*})^{*}$
                                $=(((x^{*}\rightarrow a^{*})\rightarrow(a\rightarrow a^{*}))\rightarrow a^{*})^{*}=((x\Cap a)^{*}\rightarrow a^{*})^{*}$ (by $(iG)$)
                                $=(a\rightarrow(x\Cap a))^{*}=(a\rightarrow x)^{*}$ (Prop. 2.14$(1)$).
$(4)$ It follows by $(3)$, since $x\leq_{L}a\rightarrow x$ implies $(a\rightarrow x)^{*}\leq_{L}x^{*}$.
$(5)$ From $\varphi_{a}x\leq_{L}y^{*}$, we have $y\leq_{L}(\varphi_{a}x)^{*}$. Since $\varphi_{a}$ is monotone, and applying $(4)$, we get $\varphi_{a}y\leq_{L}\varphi_{a}(\varphi_{a}x)^{*}\leq_{L}x^{*}$. The converse follows similarly.
$(6)$ Applying Proposition 2.15$(5)$, we get:
                  $\varphi_{a}\varphi_{b}a=(\varphi_{b}a)\Cap a=(a\Cap b)\Cap a=(((a\Cap b)^{*}\rightarrow a^{*})\rightarrow a^{*})^{*}$
                            $=((a\rightarrow(a\Cap b))\rightarrow a^{*})^{*}=((a\rightarrow b)\rightarrow a^{*})^{*}$
                            $=((b^{*}\rightarrow a^{*})\rightarrow a^{*})^{*}=b\Cap a=\varphi_{a}b$.
$(7)$ If $a=a\Cap b$, then $a\leq_{Q}b$, so that $a\leq_{L}b$, and by Proposition $(2)$, we have $x\Cap a=(x\Cap b)\Cap a$, for all $x\in X$. Hence $\varphi_{a}=\varphi_{a}\varphi_{b}$. ∎
Proposition 3.4.
Let $X$ be an implicative-orthomodular lattice. The following hold, for all $a,b,x,y\in X$:
$(1)$ $\varphi_{a}x=x$ iff $x\leq_{L}a;$ $(2)$ $\varphi_{a}x=0$ iff $x\leq_{L}a^{*};$
$(3)$ if $a\leq_{L}b$, then $\varphi_{a}\varphi_{b}x=\varphi_{a}x;$ $(4)$ $\varphi^{*}_{a}x=\varphi_{a}x\rightarrow y$ iff $\varphi^{*}_{a}y=\varphi_{a}y\rightarrow x;$
$(5)$ $\varphi_{a}^{2}=\varphi_{0}$ iff $\varphi_{a}1\leq_{L}\varphi_{a}^{*}1;$ $(6)$ $\varphi_{a}x\perp y$ iff $x\perp\varphi_{a}y$.
Proof.
$(1)$ If $\varphi_{a}x=x$, then $x\Cap a=x$, so that $x\leq_{Q}a$, hence $x\leq_{L}a$. Conversey, if $x\leq_{L}a$, then $x=(x\rightarrow a^{*})^{*}$ and applying Proposition 2.14$(9)$, we have: $\varphi_{a}x=x\Cap a=((x^{*}\rightarrow a^{*})\rightarrow a^{*})^{*}=(((x\rightarrow a^{*})\rightarrow a^{*})\rightarrow a^{*})^{*}=((x\Cup a^{*})\rightarrow a^{*})^{*}=(x\rightarrow a^{*})^{*}=x$.
$(2)$ Assume that $\varphi_{a}x=0$, that is $x\Cap a=0$. Hence $(x^{*}\rightarrow a^{*})\rightarrow a^{*}=1$, so that $x^{*}\rightarrow a^{*}\leq a^{*}$. Since $a^{*}\leq_{L}x^{*}\rightarrow a^{*}$, by Theorem 2.16 we get $x^{*}\rightarrow a^{*}=a^{*}$. It follows that $a=(a\rightarrow x)^{*}$, that is $a\leq_{L}x^{*}$, hence $x\leq_{L}a^{*}$. Conversely, $x\leq_{L}a^{*}$ implies $x=(x\rightarrow a)^{*}$, thus $x^{*}=x\rightarrow a=a^{*}\rightarrow x^{*}$. Hence, by $(Impl)$, $\varphi_{a}x=x\Cap a=((x^{*}\rightarrow a^{*})\rightarrow a^{*})^{*}=(((a^{*}\rightarrow x^{*})\rightarrow a^{*})\rightarrow a^{*})^{*}=(a^{*}\rightarrow a^{*})^{*}=0$.
$(3)$ If $a\leq_{L}b$, then $a^{*}=a\rightarrow b^{*}$ and we have: $\varphi_{a}\varphi_{b}x=(x\Cap b)\Cap a=(((x\Cap b)^{*}\rightarrow a^{*})\rightarrow a^{*})^{*}$. Using Proposition 2.14$(9)$, we get: $(x\Cap b)^{*}\rightarrow a^{*}=(x^{*}\Cup b^{*})\rightarrow(a\rightarrow b^{*})=a\rightarrow((x^{*}\Cup b^{*})\rightarrow b^{*})=a\rightarrow(x^{*}\rightarrow b^{*})=x^{*}\rightarrow(a\rightarrow b^{*})=x^{*}\rightarrow a^{*}$. It follows that $\varphi_{a}\varphi_{b}x=((x^{*}\rightarrow a^{*})\rightarrow a^{*})^{*}=x\Cap a=\varphi_{a}x$.
$(4)$ If $\varphi^{*}_{a}x=\varphi_{a}x\rightarrow y$, then $\varphi_{a}x\leq_{L}y^{*}$, so that $y\leq_{L}\varphi^{*}_{a}x$. Applying Propositions 3.2$(5)$, 3.3$(4)$, we get $\varphi_{a}y\leq_{L}\varphi_{a}\varphi^{*}_{a}x\leq_{L}x^{*}$, hence $\varphi^{*}_{a}y=\varphi_{a}y\rightarrow x$. The converse follows similarly.
$(5)$ If $\varphi_{a}^{2}=\varphi_{0}$, then $\varphi_{a}(\varphi_{a}1)=0$, and by $(2)$, $\varphi_{a}1\leq_{L}a^{*}=\varphi_{a}^{*}1$. Conversely, assume that $\varphi_{a}1\leq_{L}\varphi_{a}^{*}1$. Since $x\leq_{L}1$, applying Proposition 3.3$(4)$, we get $\varphi_{a}^{2}x\leq_{L}\varphi_{a}^{2}1=\varphi_{a}(\varphi_{a}1)\leq_{L}\varphi_{a}(\varphi_{a}^{*}1)\leq_{L}1^{*}=0$, so that $\varphi_{a}^{2}x=0$, for all $x\in X$. It follows that $\varphi_{a}^{2}=\varphi_{0}$.
$(6)$ It is a consequence of $(4)$. ∎
A projection is a map $f$ from a set $X$ to $X$ satisfying $f\circ f=f$ (see for example [3]). Given an implicative-orthomodular lattice $X$, by Propositions 3.2, 3.3, 3.4 a Sasaki projection $\varphi_{a}:X\longrightarrow X$ is a monotone projection from $X$ onto $[0,a]$. Moreover, the dual $\bar{\varphi}_{a}$ of $\varphi_{a}$ defined by $\bar{\varphi}_{a}x=(\varphi_{a}x^{*})^{*}=x\Cup a^{*}$ is a monoton projection from $X$ onto $[a^{*},1]$.
Definition 3.5.
Let $X$ be an implicative involutive BE algebra, and let $x,y\in X$. We say that $x$ commutes with $y$, denoted by $x\mathcal{C}y$, if $\varphi_{x}y=(x\rightarrow y^{*})^{*}$.
Lemma 3.6.
Let $X$ be an implicative involutive BE algebra. The following hold, for all $x,y\in X$:
$(1)$ $x\mathcal{C}x$, $x\mathcal{C}0$, $0\mathcal{C}x$, $x\mathcal{C}1$, $1\mathcal{C}x$, $x\mathcal{C}x^{*}$, $x^{*}\mathcal{C}x;$
$(2)$ if $x\leq_{L}y$ or $x\leq_{L}y^{*}$, then $x\mathcal{C}y;$
$(3)$ $x\mathcal{C}(y\rightarrow x)$ and $x\mathcal{C}(x^{*}\rightarrow y)$.
Proof.
$(1)$ It follows by Proposition 3.2$(1)$.
$(2)$ Since $x\leq_{L}y$ implies $x=(x\rightarrow y^{*})^{*}$ and $x\leq y$, we have $\varphi_{x}y=y\Cap x=((y^{*}\rightarrow x^{*})\rightarrow x^{*})^{*}=((x\rightarrow y)\rightarrow x^{*})^{*}=(1\rightarrow x^{*})^{*}=x=(x\rightarrow y^{*})^{*}$, that is $x\mathcal{C}y$. Similarly, from $x\leq_{L}y^{*}$ we have $x^{*}=x\rightarrow y$ and $x\leq_{L}y^{*}$, so that $x\rightarrow y^{*}=1$. It follows that $\varphi_{x}y=y\Cap x=((y^{*}\rightarrow x^{*})\rightarrow x^{*})^{*}=((x\rightarrow y)\rightarrow x^{*})^{*}=(x^{*}\rightarrow x^{*})^{*}=0=(x\rightarrow y^{*})^{*}$, hence $x\mathcal{C}y$.
$(3)$ It follows by $(2)$, since $x\leq_{L}y\rightarrow x$ and $x\leq_{L}x^{*}\rightarrow y$. ∎
Theorem 3.7.
Let $X$ be an implicative involutive BE algebra. The following are equivalent:
$(a)$ $X$ is an implicative-orthomodular lattice;
$(b)$ whenever $x\mathcal{C}y$, then $y\mathcal{C}x$.
Proof.
$(a)\Rightarrow(b)$ Suppose $X$ is an IOML, and let $x,y\in X$ such that $x\mathcal{C}y$, that is $\varphi_{x}y=y\Cap x=(x\rightarrow y^{*})^{*}$. By Proposition 2.14$(7)$, $(x\Cap y)\rightarrow(y\Cap x)=1$, so that $(x\Cap y)\rightarrow(x\rightarrow y^{*})^{*}=1$, hence $x\Cap y\leq(x\rightarrow y^{*})^{*}$. On the other hand, by $(IOM)$, $x\leq_{Q}x^{*}\rightarrow y^{*}$, thus $(x^{*}\rightarrow y^{*})\rightarrow y^{*}\leq_{Q}x\rightarrow y^{*}$. It follows that $(x\rightarrow y^{*})^{*}\leq_{L}x\Cap y$, and applying Proposition 2.14$(5)$, we get $x\Cap y=(x\rightarrow y^{*})^{*}$. Hence $\varphi_{y}x=(x\rightarrow y^{*})^{*}=(y\rightarrow x^{*})^{*}$, that is $y\mathcal{C}x$.
$(b)\Rightarrow(a)$ Let $x,y\in X$ such that $x\leq_{L}y$, that is $(x\rightarrow y^{*})^{*}=x$. By Lemma 3.6$(2)$, we have $x\mathcal{C}y$, and by $(b)$, $y\mathcal{C}x$. Hence $\varphi_{y}x=(y\rightarrow x^{*})^{*}=(x\rightarrow y^{*})^{*}=x$, so that $x\Cap y=x$, that is $x\leq_{Q}y$. Thus $x\leq_{L}y$ implies $x\leq_{Q}y$, and by Theorem 2.12, $X$ is an IOML. ∎
Corollary 3.8.
Let $X$ be an implicative involutive BE algebra. Then $X$ is an implicative-orthomodular lattice if and only if $x\mathcal{C}y$ implies $x\Cap y=y\Cap x$.
Proof.
Let $X$ be an IOML, and let $x,y\in X$ such that $x\mathcal{C}y$, that is $y\Cap x=(x\rightarrow y^{*})^{*}$. By Theorem 3.7, $x\mathcal{C}y$ implies $y\mathcal{C}x$, so that $x\Cap y=(y\rightarrow x^{*})^{*}=(x\rightarrow y^{*})^{*}=y\Cap x$. Conversely, let $X$ be an implicative involutive BE algebra, and let $x,y\in X$ such that $x\mathcal{C}y$, that is $y\Cap x=(x\rightarrow y^{*})^{*}$. By hypothesis, $y\Cap x=x\Cap y$, so that $x\Cap y=(x\rightarrow y^{*})^{*}=(y\rightarrow x^{*})^{*}$. It follows that $\varphi_{y}x=(y\rightarrow x^{*})^{*}$, hence $y\mathcal{C}x$, and by Theorem 3.7, $X$ is an IOML. ∎
Lemma 3.9.
Let $X$ be an implicative-orthomodular lattice, and let $x,y\in X$ such that $x\mathcal{C}y$. Then $x\mathcal{C}y^{*}$, $x^{*}\mathcal{C}y$ and $x^{*}\mathcal{C}y^{*}$.
Proof.
Assume that $x\mathcal{C}y$, that is $\varphi_{x}y=y\Cap x=(x\rightarrow y^{*})^{*}$. Using Proposition 2.14$(1)$, we have:
                   $\varphi_{x}y^{*}=y^{*}\Cap x=((y\rightarrow x^{*})\rightarrow x^{*})^{*}=((x\rightarrow y^{*})\rightarrow x^{*})^{*}$
                           $=((y\Cap x)^{*}\rightarrow x^{*})^{*}=(x\rightarrow(y\Cap x))^{*}$
                           $=(x\rightarrow y)^{*}=(x\rightarrow(y^{*})^{*})^{*}$.
Hence $x\mathcal{C}y^{*}$. By Theorem 3.7, we get $y^{*}\mathcal{C}x$, so that $y^{*}\mathcal{C}x^{*}$. Applying again Theorem 3.7, we have $x^{*}\mathcal{C}y^{*}$, and finally $x^{*}\mathcal{C}y$. ∎
Proposition 3.10.
Let $X$ be an implicative-orthomodular lattice, and let $x,y\in X$. Then $x\mathcal{C}y$ if and only if $(x\rightarrow y^{*})\rightarrow(x\rightarrow y)^{*}=x$.
Proof.
If $x\mathcal{C}y$, then $\varphi_{x}y=y\Cap x=(x\rightarrow y^{*})^{*}$. Using Proposition 2.14$(4)$, we have $(x\rightarrow y^{*})\rightarrow(x\rightarrow y)^{*}=(y\Cap x)^{*}\rightarrow(x\rightarrow y)^{*}=(x\rightarrow y)\rightarrow(y\Cap x)=x$. Conversely, suppose $(x\rightarrow y^{*})\rightarrow(x\rightarrow y)^{*}=x$. Then we have:
                  $x\Cap y=((x^{*}\rightarrow y^{*})\rightarrow y^{*})^{*}=((y\rightarrow x)\rightarrow y^{*})^{*}$
                           $=((y\rightarrow((x\rightarrow y^{*})\rightarrow(x\rightarrow y)^{*}))\rightarrow y^{*})^{*}$
                           $=(((x\rightarrow y^{*})\rightarrow(y\rightarrow(x\rightarrow y)^{*}))\rightarrow y^{*})^{*}$
                           $=(((x\rightarrow y^{*})\rightarrow((x\rightarrow y)\rightarrow y^{*}))\rightarrow y^{*})^{*}$
                           $=(((x\rightarrow y^{*})\rightarrow((y^{*}\rightarrow x^{*})\rightarrow y^{*}))\rightarrow y^{*})^{*}$
                           $=(((x\rightarrow y^{*})\rightarrow y^{*})\rightarrow y^{*})^{*}$ (by $(Impl)$)
                           $=((x\Cup y^{*})\rightarrow y^{*})^{*}=(y\rightarrow(x^{*}\Cap y))^{*}$
                           $=(y\rightarrow x^{*})^{*}$ (by Prop. 2.14$(1)$).
Hence $\varphi_{y}x=(y\rightarrow x^{*})^{*}$, that is $y\mathcal{C}x$, and by Theorem 3.7, $x\mathcal{C}y$. ∎
Recall that in [5] the commutativity relation on implicative-orthomodular lattices was defined based on the formula given in Proposition 3.10.
Remark 3.11.
Let $X$ be an implicative-orthomodular lattice, and let $x,y\in X$. Then $x\perp y$ implies $x\mathcal{C}y$.
Indeed, if $x\perp y$, then $y\perp x$, so that $x^{*}=x\rightarrow y$ and $y^{*}=y\rightarrow x$. It follows that $(x\rightarrow y^{*})\rightarrow(x\rightarrow y)=(x\rightarrow(y\rightarrow x))\rightarrow(x^{*})^{*}=1\rightarrow x=x$, and according to Proposition 3.10 we get $x\mathcal{C}y$.
Proposition 3.12.
Let $X$ be an implicative-orthomodular lattice, and let $x,y\in X$. Then $x\mathcal{C}y$ if and only if $x\Cap y=(x\rightarrow y^{*})^{*}$.
Proof.
Suppose $x\mathcal{C}y$, so that by Theorem 3.7, $y\mathcal{C}x$. Hence $(x\rightarrow y^{*})^{*}=(y\rightarrow x^{*})^{*}=\varphi_{y}x=x\Cap y$. Conversely, if $x\Cap y=(x\rightarrow y^{*})^{*}$, using Proposition 2.14$(4)$, we get:
                  $(y\rightarrow x^{*})\rightarrow(y\rightarrow x)^{*}=(y\rightarrow x)\rightarrow(x\rightarrow y^{*})^{*}=(y\rightarrow x)\rightarrow(x\Cap y)=y$.
By Proposition 3.10, it follows that $y\mathcal{C}x$, so that by Theorem 3.7, $x\mathcal{C}y$. ∎
Corollary 3.13.
Let $X$ be an implicative-orthomodular lattice, and let $x,y\in X$. The following are equivalent:
$(a)$ $x\mathcal{C}y;$
$(b)$ $x\Cap y=y\Cap x;$
$(c)$ $x\Cup y=y\Cup x;$
$(d)$ $\varphi_{x}y=\varphi_{y}x$.
Proof.
$(a)\Leftrightarrow(b)$ By Theorem 3.7, $x\mathcal{C}y$ iff $y\mathcal{C}x$, so that by Proposition 3.12, $x\mathcal{C}y$ iff $x\Cap y=y\Cap x$, since $(x\rightarrow y^{*})^{*}=(y\rightarrow x^{*})^{*}$.
$(a)\Leftrightarrow(c)$ By Lemma 3.9 and taking into consideration that $(a)\Leftrightarrow(b)$, we have $x\mathcal{C}y$ iff $x^{*}\mathcal{C}y^{*}$ iff $x^{*}\Cap y^{*}=y^{*}\Cap x^{*}$ iff $(x^{*}\Cap y^{*})^{*}=(y^{*}\Cap x^{*})^{*}$ iff $x\Cup y=y\Cup x$.
$(b)\Leftrightarrow(d)$ It is obvious, since $\varphi_{x}y=\varphi_{y}x$ iff $x\Cap y=y\Cap x$. ∎
In what follows, inspired from [20, Th. 5.2], we give a characterization of the commutativity relation by Sasaki projections.
Theorem 3.14.
Let $X$ be an implicative-orthomodular lattice, and let $a,b\in X$. The following are equivalent:
$(a)$ $a\mathcal{C}b;$
$(b)$ $\varphi_{a}\varphi_{b}x=\varphi_{b}\varphi_{a}x=\varphi_{a\Cap b}x$, for all $x\in X;$
$(c)$ $x\leq_{L}a$ implies $\varphi_{b}x\leq_{L}a$ and $x\leq_{L}b$ implies $\varphi_{a}x\leq_{L}b$.
Proof.
$(a)\Rightarrow(b)$ Let $a,b\in X$ such that $a\mathcal{C}b$, so that $a\Cap b=b\Cap a$. Let $x\in X$. By Proposition 2.14$(10)$, $(x\Cap b)\Cap a=(x\Cap b)\Cap(b\Cap a)$, so that $\varphi_{a}\varphi_{b}x=\varphi_{b\Cap a}\varphi_{b}x=\varphi_{a\Cap b}\varphi_{b}x$. Since $a\Cap b\leq_{L}b$, applying Proposition $(3)$, $\varphi_{a\Cap b}\varphi_{b}x=\varphi_{a\Cap b}x$. Hence $\varphi_{a}\varphi_{b}x=\varphi_{a\Cap b}x$, and similarly $\varphi_{b}\varphi_{a}x=\varphi_{a\Cap b}x$.
$(b)\Rightarrow(c)$ Taking $x:=1$ in $(b)$, we get $a\Cap b=b\Cap a$. Let $x\in X$ such that $x\leq_{L}a$, hence $x\leq_{Q}a$. It follows that $\varphi_{a}x=x\Cap a=x$, and by $(b)$, we have: $\varphi_{b}x=\varphi_{b}\varphi_{a}x=\varphi_{a\Cap b}x=\varphi_{b\Cap a}x=x\Cap(b\Cap a)\leq_{L}b\Cap a\leq_{L}a$. Similarly $x\leq_{L}b$ implies $x=\varphi_{b}x$, so that $\varphi_{a}x=\varphi_{a}\varphi_{b}x=\varphi_{a\Cap b}x=x\Cap(a\Cap b)\leq_{L}a\Cap b\leq_{L}b$.
$(c)\Rightarrow(a)$ For $x:=a$ in the first implication of $(c)$, we get $\varphi_{b}a\leq_{L}a$, hence $a\Cap b\leq_{L}a$. Since $a\Cap b\leq_{L}b$, applying Proposition 2.15$(1)$ we have $a\Cap b\leq_{L}b\Cap a$. Similarly, taking $x:=b$ in the second implication, it follows that $b\Cap a\leq_{L}b$. Since $b\Cap a\leq_{L}a$, by Proposition 2.15$(1)$ we get $b\Cap a\leq_{L}a\Cap b$. Hence $a\Cap b=b\Cap a$, so that by Corollary 3.13, $a\mathcal{C}b$. ∎
Corollary 3.15.
The composition of two Sasaki projections $\varphi_{a}$, $\varphi_{b}$ on an implicative-orthomodular lattice is commutative if and only if $a\mathcal{C}b$.
4. Implicative-Boolean center
In this section, based on the divisibility relation on implicative involutive BE algebras, we define the involutive-Boolean algebras and we give certain characterizations of these algebras. We also introduce the notion of the center $\mathcal{CC}(X)$ of an implicative-orthomodular algebra $X$, proving that $\mathcal{CC}(X)$ is an involutive-Boolean subalgebra of $X$. Moreover, we prove that an implicative involutive BE algebra $X$ is an implicative orthomodular lattice if and only if every two orthogonal elements of $X$ are contained in an implicative-Boolean subalgebra of $X$.
Definition 4.1.
Let $X$ be an implicative involutive BE algebra. We say that $x,y\in X$ satisfy the divisibility condition, denoted by $x\mathcal{D}y$, if
$(D)$ $x\rightarrow(x\rightarrow y)^{*}=x\rightarrow y^{*}$.
The notion of implicative-Boolean algebras was introduced by A. Iorgulescu in 2009 as implicative involutive BE algebras $(X,\rightarrow,^{*},1)$ satisfying condition $(D)$, for all $x,y\in X$ (see [13, Def. 3.4.12]).
Remark 4.2.
Let $X$ be an implicative involutive BE algebra, and let $x,y\in X$. Condition $(D)$ is equivalent to condition:
$(D^{\prime})$ $(y^{*}\rightarrow x)\rightarrow y=x\rightarrow y$.
Indeed, by $(D)$, $x\rightarrow y=y^{*}\rightarrow x^{*}=y^{*}\rightarrow(y^{*}\rightarrow x)^{*}=(y^{*}\rightarrow x)\rightarrow y$, that is $(D^{\prime})$. Conversely, $(D^{\prime})$ implies $x\rightarrow y^{*}=y\rightarrow x^{*}=(x\rightarrow y)\rightarrow x^{*}=x\rightarrow(x\rightarrow y)^{*}$, that is $(D)$.
Lemma 4.3.
Let $X$ be an implicative involutive BE algebra, and let $x,y\in X$. Then $x\mathcal{C}y$ if and only if $x\mathcal{D}y$.
Proof.
We have $x\mathcal{C}y$ iff $\varphi_{x}y=(x\rightarrow y^{*})^{*}$ iff $(y^{*}\rightarrow x^{*})\rightarrow x^{*}=x\rightarrow y^{*}$ iff $(x\rightarrow y)\rightarrow x^{*}=x\rightarrow y^{*}$ iff $x\rightarrow(x\rightarrow y)^{*}=x\rightarrow y^{*}$ iff $x\mathcal{D}y$. ∎
Corollary 4.4.
An implicative involutive BE algebra $X$ is an implicative-ortomodular lattice if and only if $x\mathcal{D}y$ implies $y\mathcal{D}x$.
Theorem 4.5.
Let $X$ be an implicative-orthomodular lattice. The following are equivalent:
$(a)$ $X$ is an implicative-Boolean algebra;
$(b)$ $x\Cap y=(x\rightarrow y^{*})^{*}$, for all $x,y\in X;$
$(c)$ $x\Cap y=y\Cap x$, for all $x,y\in X;$
$(d)$ $x\Cup y=y\Cup x$, for all $x,y\in X$.
Proof.
It follows by Lemma 4.3, Proposition 3.12 and Corollary 3.13. ∎
Theorem 4.6.
Let $X$ be an implicative-orthomodular lattice. The following are equivalent:
$(a)$ $X$ is an implicative-Boolean algebra;
$(b)$ $x\Cap y\leq_{L}x$, for all $x,y\in X;$
$(c)$ $x\leq_{L}x\Cup y$, for all $x,y\in X$.
Proof.
$(a)\Rightarrow(b)$ If $X$ is an implicative-Boolean algebra and $x,y\in X$, then $x\mathcal{D}y$, so that $x\mathcal{C}y$. It follows that $x\Cap y=y\Cap x\leq_{L}x$.
$(b)\Rightarrow(a)$ Let $X$ be an implicative-orthomodular lattice such that $x\Cap y\leq_{L}x$, for all $x,y\in X$. Since $x\Cap y\leq_{L}y$, applying Proposition 2.15$(1)$, we get $x\Cap y\leq_{L}y\Cap x$. Similarly, from $y\Cap x\leq_{L}x$ and $y\Cap x\leq_{L}y$, we have $y\Cap x\leq_{L}x\Cap y$, hence $x\Cap y=y\Cap x$. By Corollary 3.13, it follows that $x\mathcal{C}y$, so that, by Lemma 4.3, it follows that $X$ is an implicative-Boolean algebra.
$(a)\Rightarrow(c)$ Since $X$ is an implicative-Boolean algebra and $x,y\in X$, we have $x\mathcal{D}y$, so that $x\mathcal{C}y$. Hence $x\leq_{L}y\Cup x=x\Cup y$.
$(c)\Rightarrow(a)$ Let $X$ be an implicative-orthomodular lattice such that $x\leq_{L}x\Cup y$, for all $x,y\in X$. Since $y\leq_{L}x\Cup y$, applying Proposition 2.15$(4)$, we get $y\Cup x\leq_{L}x\Cup y$. Similarly, from $x\leq_{L}y\Cup x$ and $y\leq_{L}y\Cup x$, we have $x\Cup y\leq_{L}y\Cup x$. Hence $x\Cup y=y\Cup x$, and applying Corollary 3.13, we conclude that $X$ is an implicative-Boolean algebra. ∎
Theorem 4.7.
Let $X$ be an implicative-orthomodular lattice. The following are equivalent:
$(a)$ $X$ is an implicative-Boolean algebra;
$(b)$leq implies $\leq_{L}$.
Proof.
$(a)\Rightarrow(b)$ Let $X$ be an implicative-Boolean algebra, and let $x,y\in X$ such that $x\leq y$, that is $x\rightarrow y=1$. Since $x\rightarrow(x\rightarrow y)^{*}=x\rightarrow y^{*}$, we get $x^{*}=x\rightarrow y^{*}$, hence $x\leq_{L}y$.
$(b)\Rightarrow(a)$ By Proposition 2.14$(7)$, $(x\Cap y)\rightarrow(y\Cap x)=1$ and $(y\Cap x)\rightarrow(x\Cap y)=1$, for all $x,y\in X$. It follows that $x\Cap y\leq y\Cap x$ and $y\Cap x\leq x\Cap y$, and by $(b)$, we get $x\Cap y\leq_{L}y\Cap x$ and $y\Cap x\leq_{L}x\Cap y$. Hence $y\Cap x=x\Cap y$, for all $x,y\in X$, and applying Theorem 4.5, $X$ is an implicative-Boolean algebra. ∎
Corollary 4.8.
In any implicative-Boolean algebra, $\leq\Longleftrightarrow\leq_{L}\Longleftrightarrow\leq_{Q}$.
Definition 4.9.
The center of an implicative-orthomodular lattice $X$ is the set $\mathcal{CC}(X)=\{x\in X\mid x\mathcal{C}y$, for all $y\in X\}$.
Lemma 4.10.
Let $X$ be an implicative-orthomodular lattice. If $x,y,z\in X$ such that $x\mathcal{C}z$ and $y\mathcal{C}z$, then the following holds:
                  $x\rightarrow\ y\leq_{L}((x\rightarrow y)\rightarrow z^{*})\rightarrow((x\rightarrow y)\rightarrow z)^{*}$.
Proof.
Using Lemma 2.6$(1)$,$(8)$,$(10)$, we have successively:
                  $x^{*},y\leq_{L}x\rightarrow y;$
                  $(x\rightarrow y)\rightarrow z^{*}\leq_{L}x^{*}\rightarrow z^{*},y\rightarrow z^{*};$
                  $(x^{*}\rightarrow z^{*})^{*},(y\rightarrow z^{*})^{*}\leq_{L}((x\rightarrow y)\rightarrow z^{*})^{*};$
                  $(x^{*}\rightarrow z^{*})\rightarrow(y\rightarrow z^{*})^{*}\leq_{L}((x\rightarrow y)\rightarrow z^{*})^{*}$.
Similarly we get $(x^{*}\rightarrow z)\rightarrow(y\rightarrow z)^{*}\leq_{L}((x\rightarrow y)\rightarrow z)^{*}$.
Applying Lemma 2.6$(12)$, it follows that:
                  $((x^{*}\rightarrow z^{*})\rightarrow(y\rightarrow z^{*})^{*})^{*}\rightarrow((x^{*}\rightarrow z)\rightarrow(y\rightarrow z)^{*})$
                                                     $\leq_{L}((x\rightarrow y)\rightarrow z^{*})\rightarrow((x\rightarrow y)\rightarrow z)^{*}$,
and by Proposition 2.2$(9)$, we get:
                  $((x^{*}\rightarrow z^{*})\rightarrow(x^{*}\rightarrow z)^{*})^{*}\rightarrow((y\rightarrow z^{*})\rightarrow(y\rightarrow z)^{*})$
                                                     $\leq_{L}((x\rightarrow y)\rightarrow z^{*})\rightarrow((x\rightarrow y)\rightarrow z)^{*}$.
Since $x\mathcal{C}z$ and $y\mathcal{C}z$, by Proposition 3.10, we have $(x^{*}\rightarrow z^{*})\rightarrow(x^{*}\rightarrow z)^{*}=x^{*}$ and $(y\rightarrow z^{*})\rightarrow(y\rightarrow z)^{*}=y$. Hence $x\rightarrow\ y\leq_{L}((x\rightarrow y)\rightarrow z^{*})\rightarrow((x\rightarrow y)\rightarrow z)^{*}$. ∎
Theorem 4.11.
The center $\mathcal{CC}(X)$ of an implicative-orthomodular lattice $X$ is an implicative-Boolean subalgebra of $X$.
Proof.
According to Lemma 3.6, $0,1\in\mathcal{CC}(X)$. Let $x,y\in\mathcal{CC}(X)$, and let $z\in X$, that is $x\mathcal{C}z$ and $y\mathcal{C}z$. By Lemma 4.10, we have $x\rightarrow y\leq_{L}((x\rightarrow y)\rightarrow z^{*})\rightarrow((x\rightarrow y)\rightarrow z)^{*}$. On the other hand, by Lemma 2.6$(11)$, $((x\rightarrow y)\rightarrow z^{*})\rightarrow((x\rightarrow y)\rightarrow z)^{*}\leq_{L}x\rightarrow y$, hence $((x\rightarrow y)\rightarrow z^{*})\rightarrow((x\rightarrow y)\rightarrow z)^{*}=x\rightarrow y$. It follows that $(x\rightarrow y)\mathcal{C}z$, so that $x\rightarrow y\in\mathcal{CC}(X)$. Hence $\mathcal{CC}(X)$ is a subalgebra of $X$. Since by Lemma 4.3, $x\mathcal{C}y$ if and only if $x\mathcal{D}y$, then $\mathcal{CC}(X)$ is an implicative-Boolean algebra. ∎
$\mathcal{CC}(X)$ is called the implicative-Boolean center of $X$.
Theorem 4.12.
An implicative involutive BE algebra $X$ is an implicative orthomodular lattice if and only if every two orthogonal elements of $X$ are contained in an implicative-Boolean subalgebra of $X$.
Proof.
[…]
$\begin{array}[]{c|cccc}\rightarrow&0&x&y&1\\ \hline\cr 0&1&1&1&1\\ x&y&1&y&1\\ y&x&x&1&1\\ 1&0&x&y&1\end{array}\hskip 28.45274pt\begin{array}[]{c|cccc}\Cap&0&x&y&1\\ \hline\cr 0&0&0&0&0\\ x&0&x&0&x\\ y&0&0&y&y\\ 1&0&x&y&1\end{array}.$
Then $(Y,\rightarrow,^{*},1)$ is an implicative-Boolean subalgebra of $X$ containing $x$ and $y$.
Conversely, let $X$ be an implicative involutive BE algebra and let $x,y\in X$. Since $x$ and $y$ belong to an implicative-Boolean subalgebra of $X$, we have $x\mathcal{D}y$, hence $x\rightarrow(x\rightarrow y)^{*}=x\rightarrow y^{*}$. We also have $x\mathcal{C}y$, and applying Proposition 3.10, it follows that $x\Cup(x\rightarrow y)^{*}=(x\rightarrow(x\rightarrow y)^{*})\rightarrow(x\rightarrow y)^{*}=(x\rightarrow y^{*})\rightarrow(x\rightarrow y)^{*}=x$. Thus $X$ satisfies condition $(IOM)^{{}^{\prime\prime}}$, that is, it is an implicative-orthomodular lattice. ∎
Theorem 4.13.
$(\mathcal{SP}(\mathcal{CC}(X)),\circ,\varphi_{1})$ is an Abelian monoid.
Proof.
It is known that the composition of functions from a set into itself is associative. Since $a\mathcal{C}b$ for any $a,b\in\mathcal{CC}(X)$, then by Theorem 3.14, $\varphi_{a}\varphi_{b}=\varphi_{b}\varphi_{a}=\varphi_{a\Cap b}$, that is $\varphi_{a}\varphi_{b}\in\mathcal{SP}(\mathcal{CC}(X))$, and $\varphi_{a}$ commutes with $\varphi_{b}$. Moreover, $\varphi_{1}\in\mathcal{SP}(\mathcal{CC}(X))$, and $\varphi_{a}\varphi_{1}=\varphi_{1}\varphi_{a}=\varphi_{a}$, for any $a\in X$, so that $\varphi_{1}$ is the identity element. Hence $(\mathcal{SP}(\mathcal{CC}(X)),\circ,\varphi_{1})$ is an Abelian monoid. ∎
5. Sasaki set of projections
Similarly as in [8] for the case of orthocomplemented posets, we introduce the Sasaki sets of projections and the full Sasaki set of projections on implicative involutive BE algebras, and we investigate their properties. As a main result, we prove that an implicative involutive BE algebra is an implicative-orthomodular lattice if and only if it admits a full Sasaki set of projections.
Definition 5.1.
Let $X$ be an implicative involutive BE algebra. A subset $S$ of maps $X\longrightarrow X$ is said to be a Sasaki set of projections on $X$ if it satisfies the following conditions:
$(SS_{1})$ for each $\varphi\in S$, $x\leq_{L}y$ implies $\varphi x\leq_{L}\varphi y;$
$(SS_{2})$ if $\varphi,\varphi^{\prime}\in S$, $\varphi 1\leq_{L}\varphi^{\prime}1$ implies $\varphi\varphi^{\prime}=\varphi;$
$(SS_{3})$ for each $\varphi\in S$, $\varphi\varphi^{*}x\leq_{L}x^{*}$, for all $x\in X$.
Example 5.2.
Let $X$ be an implicative involutive BE algebra. Define $\varphi_{0},\varphi_{1}:X\longrightarrow X$ by $\varphi_{0}x=0$ and $\varphi_{1}x=x$, for all $x\in X$. The set $S=\{\varphi_{0},\varphi_{1}\}$ is a Sasaki set of projections on $X$.
Proposition 5.3.
Let $X$ be an implicative involutive BE algebra, and let $S$ be a Sasaki set of projections on $X$. The following hold, for all $\varphi,\varphi^{\prime}\in S$:
$(1)$ $\varphi 1\leq_{L}\varphi^{\prime}1$ implies $\varphi\varphi^{\prime}=\varphi=\varphi^{\prime}\varphi;$
$(2)$ $\varphi 1=\varphi^{\prime}1$ implies $\varphi=\varphi^{\prime};$
$(3)$ $\varphi^{2}=\varphi$, where $\varphi^{2}=\varphi\varphi;$
$(4)$ $\varphi 1\leq_{L}\varphi^{\prime}1$ implies $\varphi^{\prime}\varphi 1=\varphi 1;$
$(5)$ $\varphi x=0$ iff $x\leq_{L}\varphi^{*}1$.
Proof.
$(1)$ By $(SS_{3})$, $\varphi^{\prime}(\varphi^{\prime}\varphi x)^{*}\leq_{L}(\varphi x)^{*}$ and $\varphi\varphi^{\prime}(\varphi^{\prime}\varphi x)^{*}\leq_{L}\varphi(\varphi x)^{*}\leq_{L}x^{*}$. From $(SS_{2})$, $\varphi 1\leq_{L}\varphi^{\prime}1$ implies $\varphi\varphi^{\prime}=\varphi$, so that $\varphi(\varphi^{\prime}\varphi x)^{*}=\varphi\varphi^{\prime}(\varphi^{\prime}\varphi x)^{*}\leq_{L}x^{*}$. It follows that $x\leq_{L}(\varphi(\varphi^{\prime}\varphi x)^{*})^{*}$, for all $x\in X$, hence $\varphi x\leq_{L}\varphi(\varphi(\varphi^{\prime}\varphi x)^{*})^{*}$. Taking $y=(\varphi^{\prime}\varphi x)^{*}$, using $(SS_{3})$ we have $\varphi x\leq_{L}\varphi(\varphi(\varphi^{\prime}\varphi x)^{*})^{*}=\varphi(\varphi y)^{*}\leq_{L}y^{*}=((\varphi^{\prime}\varphi x)^{*})^{*}=\varphi^{\prime}\varphi x$. Thus $\varphi\leq_{L}\varphi^{\prime}\varphi$.
On the other hand, from $\varphi\varphi^{\prime}=\varphi$, we get: $(\varphi x)^{*}=(\varphi\varphi^{\prime}x)^{*}$, and applying $(SS_{3})$, we have: $\varphi^{\prime}\varphi(\varphi x)^{*}=\varphi^{\prime}\varphi(\varphi\varphi^{\prime}x)^{*}\leq_{L}\varphi^{\prime}(\varphi^{\prime}x)^{*}\leq_{L}x^{*}$. Replacing $x$ by $(\varphi x)^{*}$, it follows that $\varphi^{\prime}\varphi(\varphi(\varphi x)^{*})^{*}\leq_{L}\varphi x$. By $(SS_{3})$, $\varphi(\varphi x)^{*}\leq_{L}x^{*}$, hence $x\leq_{L}(\varphi(\varphi x)^{*})^{*}$. Since $\varphi^{\prime}\varphi$ is monotone, we get: $\varphi^{\prime}\varphi x\leq_{L}\varphi^{\prime}\varphi(\varphi(\varphi x)^{*})^{*}\leq_{L}\varphi x$. Thus $\varphi^{\prime}\varphi\leq_{L}\varphi$, and finally, $\varphi^{\prime}\varphi=\varphi$.
$(2)$, $(3)$, $(4)$ follow by $(1)$.
$(5)$ If $\varphi x=0$, by $(SS_{3})$, $\varphi(\varphi x)^{*}\leq_{L}x^{*}$, hence $\varphi(0^{*})\leq_{L}x^{*}$. It follows that $\varphi 1\leq_{L}x^{*}$, that is $x\leq_{L}(\varphi 1)^{*}$. Conversely, if $x\leq_{L}(\varphi 1)^{*}$, then by $(SS_{1})$ and $(SS_{3})$, we get $\varphi x\leq_{L}\varphi(\varphi 1)^{*}\leq_{L}1^{*}=0$, hence $\varphi x=0$. ∎
Definition 5.4.
Let $X$ be an implicative involutive BE algebra. A Sasaki set of projections $S$ on $X$ is said to be full, if for each $x\in X$, there exists $\varphi^{x}\in S$ such that $\varphi^{x}1=x$.
Remark 5.5.
By Proposition 5.3$(2)$, it follows that $\varphi^{x}$ from Definition 5.4 is unique.
Proposition 5.6.
Let $X$ be an implicative involutive BE algebra and let $S$ be a full Sasaki set of projections on $X$. Then, the following hold, for all $x,y,z\in X$ and $\varphi\in S$:
$(1)$ if $z\leq_{L}x,y$, then $z\leq\varphi^{x}(\varphi^{x}y^{*})^{*};$
$(2)$ $\varphi^{x}(\varphi^{x}y^{*})^{*}=(x\rightarrow y^{*})^{*};$
$(3)$ $\varphi^{x}x^{*}=0$.
Proof.
$(1)$ From $z\leq_{L}x$, we have $\varphi^{z}1\leq_{L}\varphi^{x}1$, and by Proposition 5.3$(1)$, we get $\varphi^{z}\varphi^{x}=\varphi^{z}$. Since $y^{*}\leq_{L}z^{*}=(\varphi^{z}1)^{*}$, then by Proposition $(5)$, we have $\varphi^{z}y^{*}=0$. It follows that $\varphi^{z}\varphi^{x}(y^{*})=\varphi^{z}(y^{*})=0$, and applying again Proposition $(5)$, we get $\varphi^{x}y^{*}\leq_{L}(\varphi^{z}1)^{*}=z^{*}$, hence $z\leq_{L}(\varphi^{x}y^{*})^{*}$. From $\varphi^{z}1=z\leq_{L}x=\varphi^{x}1$, by Proposition 5.3$(4)$ we have $\varphi^{x}\varphi^{z}1=\varphi^{z}1$, thus $\varphi^{x}z=z$. Finally, $z=\varphi^{x}z\leq_{L}\varphi^{x}(\varphi^{x}y^{*})^{*}$.
$(2)$ Since $(\varphi^{x}y^{*})^{*}\leq_{L}1$, we have $\varphi^{x}(\varphi^{x}y^{*})^{*}\leq_{L}\varphi^{x}1=x$. Moreover, by $(SS_{3})$, $\varphi^{x}(\varphi^{x}y^{*})^{*}\leq_{L}(y^{*})^{*}=y$. Applying Lemma 2.2$(8)$, we get $\varphi^{x}(\varphi^{x}y^{*})^{*}\leq_{L}(x\rightarrow y^{*})^{*}$.
On the other hand, from $(x\rightarrow y^{*})^{*}\leq_{L}x,y$ (Lemma 2.6$(4)$), using $(1)$, $(x\rightarrow y^{*})^{*}\leq_{L}\varphi^{x}(\varphi^{x}y^{*})^{*}$. We conclude that $\varphi^{x}(\varphi^{x}y^{*})^{*}=(x\rightarrow y^{*})^{*}$.
$(3)$ Since $x^{*}\leq_{L}x^{*}$ implies $x^{*}\leq_{L}(\varphi^{x}1)^{*}$, using Proposition $(5)$, we get $\varphi^{x}x^{*}=0$. ∎
Proposition 5.7.
Let $X$ be an implicative involutive BE algebra, let $S$ be a Sasaki set of projections on $X$, and let $\varphi\in S$. The following holds, for all $x,y\in X$:
                                            $\varphi(x\rightarrow y)=\varphi^{*}x^{*}\rightarrow\varphi y$.
Proof.
Since $X$ is an IOML, we have $\leq_{Q}\iff\leq_{L}$. Denote $z=\varphi^{*}x^{*}\rightarrow\varphi y$. From $\varphi x^{*}\leq_{L}\varphi^{*}x^{*}\rightarrow\varphi y=z$, we get $z^{*}\leq_{L}\varphi^{*}x^{*}$, and by $(SS_{3})$, $\varphi z^{*}\leq_{L}\varphi\varphi^{*}x^{*}\leq_{L}x$. Hence $x^{*}\leq_{L}\varphi^{*}z^{*}$. Similarly, from $\varphi y\leq_{L}\varphi^{*}x^{*}\rightarrow\varphi y=z$, we have $y\leq_{L}\varphi^{*}z^{*}$. By Lemma 2.6$(10)$, $x\rightarrow y\leq_{L}\varphi^{*}z^{*}$, and by $(SS_{3})$, $\varphi(x\rightarrow y)\leq_{L}\varphi\varphi^{*}z^{*}\leq_{L}z=\varphi^{*}x^{*}\rightarrow\varphi y$. Moreover, from $x^{*},y\leq_{L}x\rightarrow y$, we have $\varphi x^{*},\varphi y\leq_{L}\varphi(x\rightarrow y)$. Applying again Lemma 2.6$(10)$, we get $\varphi^{*}x^{*}\rightarrow\varphi y\leq_{L}\varphi(x\rightarrow y)$. We conclude that $\varphi(x\rightarrow y)=\varphi^{*}x^{*}\rightarrow\varphi y$. ∎
Proposition 5.8.
Let $X$ be an implicative involutive BE algebra, let $S$ be a full Sasaki set of projections on $X$ and let $\varphi\in S$. Then $\varphi^{x}y=y\Cap x$, for all $x,y\in X$.
Proof.
For all $x,y\in X$, we have:
                           $y\Cap x=((y^{*}\rightarrow x^{*})\rightarrow x^{*})^{*}=(x\rightarrow(y^{*}\rightarrow x^{*})^{*})^{*}$
                                    $=\varphi^{x}(\varphi^{x}(y^{*}\rightarrow x^{*})^{*})^{*}$ (Prop. 5.6(2))(2))( 2 ) )
                                    $=\varphi^{x}(\varphi^{x}(x\rightarrow y)^{*})^{*}$
                                    $=\varphi^{x}(\varphi^{x}\varphi^{x}(\varphi^{x}y)^{*})^{*}$ (Prop. 5.6(2))(2))( 2 ) )
                                    $=\varphi^{x}(\varphi^{x}(\varphi^{x}y)^{*})^{*}$ (Prop. 5.$(3))$
                                    $=\varphi^{x}(x\rightarrow y)$ (Prop. 5.6(2))(2))( 2 ) )
                                    $=\varphi^{x}(y^{*}\rightarrow x^{*})$
                                    $=(\varphi^{x}y)^{*}\rightarrow\varphi^{x}(x^{*})$ (Prop. 5.7)
                                    $=(\varphi^{x}y)^{*}\rightarrow 0$ (Prop. 5.6(3))(3))( 3 ) )
                                    $=\varphi^{x}y$. ∎
Based on Sasaki set of projections, we give a characterization theorem for implicative-orthomodular lattices.
Theorem 5.9.
An implicative involutive BE algebra is an implicative-orthomodular lattice if and only if it admits a full Sasaki set of projections.
Proof.
Let $(X,\rightarrow,^{*},1)$ be an implicative involutive BE algebra, and let $S$ be a full Sasaki set of projections on $X$. Consider $x,y\in X$ such that $x\leq y$ and $y\leq_{L}x$. Then there exists $\varphi^{x},\varphi^{y}\in S$ such that $\varphi^{x}1=x$ and $\varphi^{y}1=y$. Applying Proposition 5.6$(2)$, it follows that $\varphi^{x}(\varphi^{x}y)^{*}=(x\rightarrow y)^{*}$. From $y\leq_{L}x$, we have $\varphi^{y}1\leq_{L}\varphi^{x}1$, and by Proposition 5.3$(1)$, $\varphi^{x}\varphi^{y}1=\varphi^{y}1$, that is $\varphi^{x}y=y$. Then $(\varphi^{x}y)^{*}=y^{*}$, so that $\varphi^{x}(\varphi^{x}y)^{*}=\varphi^{x}y^{*}$, hence $\varphi^{x}y^{*}=(x\rightarrow y)^{*}=0$, since $x\leq y$. By Proposition $(5)$, we get $y^{*}\leq_{L}(\varphi^{x}1)^{*}=x^{*}$, that is $x\leq_{L}y$. Thus $x=y$, so that, according to Theorem 2.16, $X$ is an implicative-orthomodular lattice.
Conversely, if $X$ is an IOML, let $S=\mathcal{SP}(X)=\{\varphi_{a}\mid a\in X\}$, the set of all Sasaki projections on $X$. By Propositions 3.2$(5)$, 3.3$(4)$, it follows that $(SS_{1})$ and $(SS_{3})$ hold for any $\varphi_{a}\in S$. If $\varphi_{a},\varphi_{b}\in S$ such that $\varphi_{a}1\leq_{L}\varphi_{b}1$, then $a\leq_{L}b$, and applying Proposition $(2)$, we have $(x\Cap b)\Cap a=x\Cap a$, for all $x\in X$. Hence $\varphi_{a}\varphi_{b}=\varphi_{a}$, so that condition $(SS_{2})$ is verified. It follows that $S$ is a Sasaki set of projections on $X$. For any $x\in X$, taking $\varphi^{x}:=\varphi_{x}\in S$, we have $\varphi^{x}1=\varphi_{x}1=x$, hence $S$ is a full Sasaki set of projections on $X$. ∎
6. Connection with Dacey and Sasaki spaces
Sasaki spaces were introduced in [16], based on Sasaki maps on orthogonality spaces. In this section, we investigate these notions in the case of implicative involutive BE algebras, and we define and characterize the Dacey spaces. We prove that the associated orthogonality space of a complete implicative-orthomodular lattice is a Sasaki space as well as a Dacey space. Finally, we show that, in the case when an implicative involutive BE algebra admits a full Sasaki set of projections, its associated orthogonality space is a Sasaki space.
Let $(X,\rightarrow,^{*},1)$ be an implicative involutive BE algebra, and let $(X^{\prime},\perp)$ be its associated orthogonality space, where $X^{\prime}=X\setminus\{0\}$. If $A,B\subseteq X^{\prime}$, denote: $A^{\perp}=\{x\in X\mid x\perp a$, for all $a\in A\}$, $A\wedge B=A\cap B$, $A\vee B=(A^{\perp}\wedge B^{\perp})^{\perp}$. Clearly, $A\subseteq B$ implies $B^{\perp}\subseteq A^{\perp}$.
Remark 6.1.
For any $A\subseteq X^{\prime}$, we have $A\subseteq A^{\perp\perp}$ and $A^{\perp\perp\perp}=A^{\perp}$.
Indeed, let $x\in A$. For any $y\in A^{\perp}$, we have $y\perp A$, that is $x\perp y$. Hence $x\in A^{\perp\perp}$, so that $A\subseteq A^{\perp\perp}$. It follows that $A^{\perp\perp\perp}\subseteq A^{\perp}$, and $A\subseteq A^{\perp\perp}$ implies also $A^{\perp}\subseteq(A^{\perp})^{\perp\perp}=A^{\perp\perp\perp}$. Hence $A^{\perp\perp\perp}=A^{\perp}$.
A subset of mutually orthogonal elements of an orthogonality space is called a $\perp$-set. The subset $A$ of $X^{\prime}$ is called orthoclosed if $A=A^{\perp\perp}$, and denote by $\mathcal{C}(X^{\prime},\perp)$ the set of all orthoclosed subsets of $X^{\prime}$. For any $A,B\in\mathcal{C}(X^{\prime},\perp)$, define $A\leq_{L}B$ if $A\subseteq B$ and $A\rightarrow B=(A\cap B^{\perp})^{\perp}$. We can see that $A^{*}=A\rightarrow\emptyset=(A\cap\emptyset^{\perp})^{\perp}=A^{\perp}$.
Remark 6.2.
For any $A,B\in\mathcal{C}(X^{\prime},\perp)$, we have $A^{\perp},A\cap B,A\vee B\in\mathcal{C}(X^{\prime},\perp)$.
Indeed, $A^{\perp}=A^{\perp\perp\perp}$ implies $A^{\perp}\in\mathcal{C}(X^{\prime},\perp)$. Since $A\cap B\subseteq A,B$ implies $(A\cap B)^{\perp\perp}\subseteq A,B$, we get $(A\cap B)^{\perp\perp}\subseteq A\cap B$. On the other hand $A\cap B\subseteq(A\cap B)^{\perp\perp}$, so that $(A\cap B)^{\perp\perp}=A\cap B$, hence $A\cap B\in\mathcal{C}(X^{\prime},\perp)$. Since $A\vee B=(A^{\perp}\cap B^{\perp})^{\perp}$, it follows that $A\vee B\in\mathcal{C}(X^{\prime},\perp)$.
Lemma 6.3.
Let $(X,\rightarrow,^{*},1)$ be an implicative involutive BE algebra, and let $(X^{\prime},\perp)$ be its associated orthogonality space. Then $(\mathcal{C}(X^{\prime},\perp),\rightarrow,^{*},X^{\prime})$ is an implicative involutive BE algebra.
Proof.
Obviously, $\mathcal{C}(X^{\prime},\perp)$ is involutive and it verifies the axioms of BE algebras. Moreover, $(A\rightarrow B)\rightarrow A=((A\cap B^{\perp})^{\perp}\cap A^{\perp})^{\perp}=((A^{\perp}\vee B)\cap A^{\perp})^{\perp}=(A^{\perp})^{\perp}=A$, hence axiom $(Impl)$ is also verified. ∎
J. Paseka and T. Vetterlein defined in [19] the normal orthogonality spaces and characterized their category. A decomposition of an orthogonality space $(X,\perp)$ is a pair $(A_{1},A_{2})$ of non-empty subsets of $X$ such that $A_{1}=A_{2}^{\perp}$ and $A_{2}=A_{1}^{\perp}$. Clearly, $A_{1}\perp A_{2}$ and $(A_{1}\cup A_{2})^{\perp}=\emptyset$. A maximal $\perp$-set of $X$ is called a block, and a block partition is a partition of a block of $X$ into non-empty subsets called cells.
Definition 6.4.
([19]) An orthogonality space $(X,\perp)$ is called normal if, for any block partition $(E_{1},E_{2})$ there is a unique decomposition $(\overline{E_{1}},\overline{E_{2}})$ such that $E_{1}\subseteq\overline{E_{1}}$, $E_{2}\subseteq\overline{E_{2}}$.
Lemma 6.5.
([19, Prop. 2.5]) Let $(X,\perp)$ be an orthogonality space. Then $X$ is normal if and only if for any block partition $(E_{1},E_{2})$, the pair $(E_{2}^{\perp},E_{1}^{\perp})$ is a decomposition of $X$.
We prove for implicative involutive BE algebras a result from [19, Prop. 2.8].
Proposition 6.6.
Let $(X,\rightarrow,^{*},1)$ be an implicative involutive BE algebra such that the associated orthogonality space $(X^{\prime},\perp)$ is normal. For any maximal $\perp$-subset $E$ of $X$, denote $\mathcal{B}_{E}=\{A^{\perp\perp}\mid A\subseteq E\}$. Then $(\mathcal{B}_{E},\rightarrow,^{\perp},X)$ is an implicative-Boolean subalgebra of $\mathcal{C}(X^{\prime},\perp)$.
Proof.
Clearly, $\mathcal{B}_{E}\subseteq\mathcal{C}(X^{\prime},\perp)$ and $\emptyset\in\mathcal{B}_{E}$. Obviously, $E^{\perp\perp}\subseteq X$. Assume that there exists $x\in X$, $x\notin E^{\perp\perp}$. Then $E^{\perp}=\emptyset$ (indeed, if $e\in E^{\perp}$, then $e\perp E$, so that $E\cup\{e\}$ is a $\perp$-set contained in $X$, hence $E$ is not maximal, a contradiction). It follows that $E^{\perp\perp}=X$, with $E\subseteq E$, that is $X\in\mathcal{B}_{E}$. One can easy prove that $\mathcal{B}_{E}$ is closed under arbitrary meets and joins (see also [19, Prop. 2.8]). Let $A^{\perp\perp}\in\mathcal{B}_{E}$. Since $(A,E\setminus A)$ is a block partition, then by Lemma 6.5, $((E\setminus A)^{\perp},A^{\perp})$ is a decomposition of $X$. Hence $A^{\perp}\perp(E\setminus A)^{\perp}$, so that $A^{\perp}=(E\setminus A)^{\perp\perp}$ and $(A^{\perp\perp})^{\perp}=A^{\perp}=(E\setminus A)^{\perp\perp}\in\mathcal{B}_{E}$, hence $\mathcal{B}_{E}$ is closed under . For any $A^{\perp\perp},B^{\perp\perp}\in\mathcal{B}_{E}$, we have $(A^{\perp\perp}\rightarrow B^{\perp\perp})^{\perp\perp}=(A^{\perp\perp}\cap B^{\perp})^{\perp}\in\mathcal{B}_{E}$. Moreover, $A^{\perp\perp}\rightarrow(A^{\perp\perp}\rightarrow B^{\perp\perp})^{\perp}=A^{\perp}\vee(A^{\perp}\vee B^{\perp\perp})^{\perp}=A^{\perp}\vee(A^{\perp\perp}\cap B^{\perp})=A^{\perp}\vee B^{\perp}=(A^{\perp\perp}\cap B^{\perp\perp})^{\perp}=A^{\perp\perp}\rightarrow(B^{\perp\perp})^{\perp}$, that is $\mathcal{B}_{E}$ satisfies condition $\mathcal{D}$. Hence $\mathcal{B}_{E}$ is an implicative-Boolean subalgebra of $\mathcal{C}(X^{\prime},\perp)$. ∎
Definition 6.7.
Let $(X,\rightarrow,^{*},1)$ be an implicative involutive BE algebra. The orthogonality space $(X^{\prime},\perp)$ is a Dacey space if $\mathcal{C}(X^{\prime},\perp)$ is an implicative-orthomodular lattice.
Proposition 6.8.
Let $(X,\rightarrow,^{*},1)$ be an implicative involutive BE algebra, and let $(X^{\prime},\perp)$ be its associated orthogonality space. The following are equivalent:
$(a)$ $(X^{\prime},\perp)$ is a Dacey space;
$(b)$ Every two orthogonal elements of $\mathcal{C}(X^{\prime},\perp)$ are contained in an involutive-Boolean subalgebra of $\mathcal{C}(X^{\prime},\perp)$.
Proof.
It follows by Theorem 4.12. ∎
Definition 6.9.
Let $(X^{\prime},\perp)$ be the associated orthogonality space of an implicative involutive BE algebra, and let $A$ be an orthoclosed subset of $X^{\prime}$. A Sasaki map to $A$ is a map $\varphi:\complement A^{\perp}\longrightarrow A$ satisfying the following conditions:
$(SM_{1})$ $\varphi x=x$, for all $x\in A;$
$(SM_{2})$ for any $x,y\in\complement A^{\perp}$, $\varphi x\perp y$ if and only if $x\perp\varphi y$.
We call $(X^{\prime},\perp)$ a Sasaki space if for any orthoclosed subset $A$ of $X^{\prime}$, there exists a Sasaki map $\varphi$ to $A$.
According to [16], any Sasaki space is a Dacey space.
We say that an implicative involutive BE algebra (implicative-orthomodular lattice) $(X,\rightarrow,^{*},0,1)$ is complete if the corresponding lattice $(X,\wedge,\vee,0,1)$ (see Remark 2.11) is complete.
For any $x\in X$, denote $\downarrow x=\{y\in X\mid y\leq_{L}x\}$.
Lemma 6.10.
Let $X$ be an implicative involutive BE algebra. The following hold, for any $x,y\in X$:
$(1)$ $(\downarrow x)^{\perp}\cap X^{\prime}=\downarrow x^{*}\cap X^{\prime}=x^{\perp}\cap X^{\prime};$
$(2)$ $(\downarrow x)\cap(\downarrow y)=\downarrow(x\rightarrow y^{*})^{*};$
$(3)$ $\downarrow(x\rightarrow y)=(\downarrow x)\rightarrow(\downarrow y)$.
If $X$ is complete and $Y\subseteq X$ is non-empty, then:
$(4)$ $\bigwedge_{y\in Y}(\downarrow y)=\downarrow(\bigwedge_{y\in Y}y);$
$(5)$ $Y^{\perp}=\downarrow(\bigwedge_{y\in Y}y^{*})$.
Proof.
$(1)$ First we show that $\downarrow x\cap\downarrow x^{*}=\{0\}$ and $\downarrow x\cap x^{\perp}=\{0\}$. Indeed, $u\in\downarrow x\cap\downarrow x^{*}$ implies $u\leq_{L}x$ and $u\leq_{L}x^{*}$, that is $u\leq_{L}x$ and $x\leq_{L}u^{*}$. It follows that $u\leq_{L}u^{*}$, so that $u^{*}=u\rightarrow u=1$, hence $u=0$. Similarly, for any $u\in\downarrow x\cap x^{\perp}$ we have $u\leq_{L}x$ and $u\perp x$, thus $u\leq_{L}x$ and $u\leq_{L}x^{*}$. Finally, we get $u\leq_{L}x\leq_{L}u^{*}$, hence $u=0$.
For any $u\in(\downarrow x)^{\perp}$, we have $u\perp x$ iff $u^{*}=u\rightarrow x$ iff $u\leq_{L}x^{*}$ iff $u\in\downarrow x^{*}$, hence $(\downarrow x)^{\perp}\subseteq\downarrow x^{*}$. Conversely, if $u\in\downarrow x^{*}$, then $u\leq_{L}x^{*}$. For all $v\in X^{\prime}$, $v\leq_{L}x$, we have $x^{*}\leq_{L}v^{*}$. It follows that $u\leq_{L}v^{*}$, so that $u^{*}=u\rightarrow v$, that is $u\perp v$. Hence $u\in(\downarrow x)^{\perp}$, thus $\downarrow x^{*}\subseteq(\downarrow x)^{\perp}$. It follows that $(\downarrow x)^{\perp}\cap X^{\prime}=\downarrow x^{*}\cap X^{\prime}$.
Moreover, $u\in\downarrow x^{*}$ iff $u\leq_{L}x^{*}$ iff $u^{*}=u\rightarrow x$ iff $u\perp x$ iff $x\in x^{\perp}$, so that $\downarrow x^{*}\cap X^{\prime}=x^{\perp}\cap X^{\prime}$.
$(2)$ If $u\in(\downarrow x)\cap(\downarrow y)$, we have $u\leq_{L}x,y$, hence $u\leq_{L}(x\rightarrow y^{*})^{*}$. It follows that $u\in\downarrow(x\rightarrow y^{*})^{*}$, so that $(\downarrow x)\cap(\downarrow y)\subseteq\downarrow(x\rightarrow y^{*})^{*}$. Conversely, $\downarrow(x\rightarrow y^{*})^{*}$ implies $u\leq_{L}(x\rightarrow y^{*})^{*}\leq_{L}x,y$, that is $u\in(\downarrow x)\cap(\downarrow y)$, hence $\downarrow(x\rightarrow y^{*})^{*}\subseteq(\downarrow x)\cap(\downarrow y)$.
$(3)$ Using $(2)$, we have:
                  $\downarrow(x\rightarrow y)=\downarrow(x\rightarrow y)^{**}=(\downarrow(x\rightarrow y)^{*})^{\perp}=(\downarrow x\cap\downarrow y^{*})^{\perp}$
                                 $=(\downarrow x\cap(\downarrow y)^{\perp})^{\perp}=(\downarrow x)\rightarrow(\downarrow y)$.
$(4)$ For any $u\in X^{\prime}$, we have: $u\in\bigwedge_{y\in Y}(\downarrow y)$ iff $u\in\downarrow y$, for all $y\in Y$ iff $u\leq_{L}y$, for all $y\in Y$ iff $u\leq_{L}\bigwedge_{y\in Y}y$, for all $y\in Y$ iff $u\in\downarrow(\bigwedge_{y\in Y}y)$. Hence $\bigwedge_{y\in Y}(\downarrow y)=\downarrow(\bigwedge_{y\in Y}y)$.
$(5)$ Using $(3)$, we get:
                  $Y^{\perp}=\{x\in X\mid x\perp y$, for all $y\in Y\}=\{x\in X\mid x\leq_{L}y^{*}$, for all $y\in Y\}$
                        =$=\{x\in X\mid x\in\downarrow y^{*}$, for all $y\in Y\}=\bigwedge_{y\in Y}(\downarrow y^{*})=\downarrow(\bigwedge_{y\in Y}y^{*})$. ∎
Following the idea from [16, Prop. 2.8], in the next result we show that any Sasaki projection on a complete implicative-orthomodular lattice induces a Sasaki map on the associated orthogonality space.
Proposition 6.11.
Let $(X,\rightarrow,^{*},1)$ be a complete implicative-orthomodular lattice. Then the associated orthogonality space $(X^{\prime},\perp)$ is a Sasaki space with $\mathcal{C}(X^{\prime},\perp)=\{\downarrow x\cap X^{\prime}\mid x\in X\}$, where $\downarrow x=\{y\in X\mid y\leq_{L}x\}$. For any $A=\downarrow x\cap X^{\prime}\in\mathcal{C}(X^{\prime},\perp)$, the restriction and corestriction of the Sasaki projection $\varphi_{x}:X\longrightarrow X$ to a map $\varphi:\complement A^{\perp}\longrightarrow A$ is a Sasaki map to $A$.
Proof.
The proof follows similar to that in [16, Prop. 2.8]. By Lemma 6.10, for any non-empty subset $Y\subseteq X$ we have $Y^{\perp}=\downarrow(\bigwedge_{y\in Y}y^{*})$. Since for any $A\in\mathcal{C}(X^{\prime},\perp)$, $A=Y^{\perp}$ for some $Y\subseteq X$, it follows that $\mathcal{C}(X^{\prime},\perp)=\{\downarrow x\cap X^{\prime}\mid x\in X\}$. Let $A\in\mathcal{C}(X^{\prime},\perp)$ with $A=\downarrow x\cap X^{\prime}$ for some $x\in X$. Consider the map $\varphi_{A}:\complement A^{\perp}\longrightarrow A$ defined as the restriction and corestriction of the Sasaki projection $\varphi_{x}:X\longrightarrow X$ ($\varphi_{A}=\varphi{{}_{x}}_{\mid\complement A^{\perp}}$). We have:
                  $\complement A^{\perp}=\complement(\downarrow x\cap X^{\prime})^{\perp}=\complement(\phi x)^{\perp}=\complement\psi(x^{*})=X^{\prime}\setminus(\downarrow x^{*}\cap X^{\prime})=X^{\prime}\setminus\downarrow x^{*}$.
It follows that, for any $y\in\complement A^{\perp}$, $y\notin\downarrow x^{*}$, so that $y\nleq x^{*}$. Applying Proposition 3.4$(2)$, $\varphi_{x}y\neq 0$, hence $\varphi_{A}$ is well defined. For any $y\in A$, we have $y\in\downarrow x$, that is $y\leq x$. By Proposition 3.4$(1)$ we get $\varphi_{x}y=x$, so that condition $(SM_{1})$ is verified. Since for all $x,y\in\complement A^{\perp}$, by Proposition 3.4$(6)$ we have $\varphi x\perp y$ iff $x\perp\varphi y$, condition $(SM_{2})$ is also satisfied. Hence $\varphi_{A}$ is a Sasaki map. ∎
Remark 6.12.
If $(X,\rightarrow,^{*},1)$ be a complete implicative-orthomodular lattice, then $X$ and $\mathcal{C}(X^{\prime},\perp)$ are isomorphic as implicative-orthomodular lattice.
Indeed, define $\phi:X\longrightarrow\mathcal{C}(X^{\prime},\perp)$ by $\phi x=\downarrow x\cap X^{\prime}$. Obviously, $x\neq y$ implies $\downarrow x\cap X^{\prime}\neq\downarrow y\cap X^{\prime}$. Since $\mathcal{C}(X^{\prime},\perp)=\{\downarrow x\cap X^{\prime}\mid x\in X\}$, for any $A\in\mathcal{C}(X^{\prime},\perp)$ there exists $x\in X$ such that $\phi x=A$. Applying again Lemma 6.10, we get:
                  $\phi(x\rightarrow y)=\downarrow(x\rightarrow y)\cap X^{\prime}=(\downarrow x\rightarrow\downarrow y)\cap X^{\prime}=\phi x\rightarrow\phi y$,
                  $\phi(x^{*})=\downarrow x^{*}\cap X^{\prime}=(\downarrow x)^{\perp}\cap X^{\prime}=(\phi x)^{\perp}$.
It folllows that $\phi$ is an isomorphism of implicative-orthomodular lattices.
Proposition 6.13.
If a complete implicative involutive BE algebra admits a full Sasaki set of projections, then its associated orthogonality space is a Sasaki space.
Proof.
Let $(X,\rightarrow,^{*},1)$ be a complete implicative involutive BE algebra, and let $(X^{\prime},\perp)$ be its associated orthogonality space ($X^{\prime}=X\setminus\{0\}$). Since $X$ admits a full Sasaki set of projections, then by Theorem 5.9 it is a complete implicative-orthomodular lattice. Applying Proposition 6.11, for any orthoclosed subset $A$ of $X^{\prime}$ there exists a Sasaki map to $A$. Hence $(X^{\prime},\perp)$ is a Sasaki space. ∎
Example 6.14.
[…]
Then $(X,\rightarrow,^{*},1)$, where $x^{*}=x\rightarrow 0$ is an implicative involutive BE algebra, but it is not an implicative-orthomodular lattice. Indeed, $a\Cap(d\rightarrow a)=b\neq a$, so that condition $(IOM)$ is not satisfied. Denoting $X^{\prime}=X\setminus\{0\}$, then $(X^{\prime},\perp)$ is an orthogonality space, with $\perp$ defined in Section 1. We can see that $a^{\perp}=\{c,d\}$, $b^{\perp}=\{d\}$, $c^{\perp}=\{a\}$, $d^{\perp}=\{a,b\}$. Denote $A=\{a\}$, $B=\{d\}$, $C=\{a,b\}$, $D=\{c,d\}$, hence $\mathcal{C}(X^{\prime},\perp)=\{\emptyset,A,B,C,D,X^{\prime}\}$.
We have $C^{\perp}=\{d\}$ and $\complement C^{\perp}=\{a,b,c,1\}$. Assume that there exists a Sasaki map $\varphi:\complement C^{\perp}\longrightarrow C$, hence by $(SS_{1})$, $\varphi a=a$ and $\varphi b=b$. For $a,c\in\complement C^{\perp}$ we have $\varphi a=a\perp c$, but $a\not\perp\varphi c\in\{a,b\}$. Hence condition $(SS_{2})$ is not satisfied, so that there are no Sasaki maps on $C=\{a,b\}$. Thus $(X^{\prime},\perp)$ is not a Sasaki space.
[…]
$\begin{array}[]{c|cccccc}\rightarrow&\emptyset&A&B&C&D&X^{\prime}\\ \hline\cr\emptyset&X^{\prime}&X^{\prime}&X^{\prime}&X^{\prime}&X^{\prime}&X^{\prime}\\ A&D&X^{\prime}&D&X^{\prime}&D&X^{\prime}\\ B&C&C&X^{\prime}&C&X^{\prime}&X^{\prime}\\ C&B&X^{\prime}&B&X^{\prime}&D&X^{\prime}\\ D&A&A&X^{\prime}&C&X^{\prime}&X^{\prime}\\ X^{\prime}&\emptyset&A&B&C&D&X^{\prime}\end{array}\hskip 28.45274pt\begin{array}[]{c|ccccccc}\Cap&\emptyset&A&B&C&D&X^{\prime}\\ \hline\cr\emptyset&\emptyset&\emptyset&\emptyset&\emptyset&\emptyset&\emptyset\\ A&\emptyset&A&\emptyset&C&\emptyset&A\\ B&\emptyset&\emptyset&B&\emptyset&D&B\\ C&\emptyset&A&\emptyset&C&\emptyset&C\\ D&\emptyset&\emptyset&B&\emptyset&D&D\\ X^{\prime}&\emptyset&A&B&C&D&X^{\prime}\end{array}.$
$(\mathcal{C}(X^{\prime},\perp),\rightarrow,^{*},X^{\prime})$ is an implicative involutive BE algebra, but it is not an IOML. Indeed, since $A\Cap(B\rightarrow A)=C\neq A$, condition $(IOM)$ is not satisfied. It follows that $(X^{\prime},\perp)$ is not a Dacey space.
Example 6.15.
[…]
Then $(X,\rightarrow,^{*},1)$, where $x^{*}=x\rightarrow 0$ is an implicative-orthomodular lattice. Denoting $X^{\prime}=X\setminus\{0\}$, then $(X^{\prime},\perp)$ is an orthogonality space, with $\perp$ defined in Section 1. We can see that $a^{\perp}=\{d\}$, $b^{\perp}=\{c\}$, $c^{\perp}=\{b\}$, $d^{\perp}=\{a\}$ and $\mathcal{C}(X^{\prime},\perp)=\{\emptyset,\{a\},\{b\},\{c\},\{d\},X^{\prime}\}$. The maps:
                                    $\varphi^{a}:\complement a^{\perp}\longrightarrow\{a\}$, $\varphi^{a}x=a$,
                                    $\varphi^{b}:\complement b^{\perp}\longrightarrow\{b\}$, $\varphi^{b}x=b$,
                                    $\varphi^{c}:\complement c^{\perp}\longrightarrow\{c\}$, $\varphi^{c}x=c$,
                                    $\varphi^{d}:\complement d^{\perp}\longrightarrow\{d\}$, $\varphi^{d}x=d$,
are Sasaki maps, hence $(X^{\prime},\perp)$ is a Sasaki space. It follows that $(X^{\prime},\perp)$ is also a Dacey space.
Example 6.16.
Consider the orthogonality space $(X^{\prime},\perp)$ from Example 6.14, where $\{a,d\}$, $\{a,c\}$, $\{b,d\}$ are the blocks of $X^{\prime}$. Denoting $E_{1}=\{a\}$, $E_{2}=\{c\}$, $\overline{E_{1}}=\{a\}$, $\overline{E_{2}}=\{c,d\}$, we have $E_{1}\subseteq\overline{E_{1}}$, $E_{2}\subseteq\overline{E_{2}}$, $\overline{E_{1}}^{\perp}=a^{\perp}=\{c,d\}=\overline{E_{2}}$ and $\overline{E_{2}}^{\perp}=\{c,d\}^{\perp}=\{a\}=\overline{E_{1}}$. It follows that $(\overline{E_{1}},\overline{E_{2}})$ is the unique decomposition of the block partition $(E_{1},E_{2})$. Similarly, for the blocks $\{a,d\}$, $\{b,d\}$ there are unique decompositions satisfying the condition from Definition 6.4. Hence $(X^{\prime},\perp)$ is a normal orthogonality space.
Declarations
Ethical Approval
This declaration is not applicable.
Competing interests
The author has no relevant financial or non-financial interests to disclose.
Funding
No funds, grants, or other support was received.
Availability of data and materials
This declaration is not applicable.
Authors’ contributions
This declaration is not applicable.